### v1

# Controlling superradiant phase transition in quantum Rabi model

Abstract
In the ultrastrong-coupling regime, the quantum Rabi model can exhibit quantum phase transition (QPT) when the ratio of the qubit transition frequency to the frequency of the cavity field approaches infinity. However, it is challenging to control the QPT in few-body systems because of the limited coupling strength and the $A^{2}$ terms. Here, we propose a practical scheme to manipulate the QPT of quantum Rabi model in the strong-coupling regime. By applying a periodic frequency modulation to the two-level system in a standard quantum Rabi model in the strong-coupling regime, an anisotropic quantum Rabi model with ultrastrong and tunable coupling strengths for rotating and counter-rotating terms is obtained. The ground-state and excitation energy of this model in terms of the modulation parameters are studied. We find that the QPT of quantum Rabi model can be observed in the strong-coupling regime and externally controlled by the modulation.
I Introduction
Quantum phase transition (QPT) QPT1999 has long been an important and fundamental phenomenon in many-body systems, and has attracted remarkable amount of study interest due to its special features. The QPT occurs when the non-thermal parameter, such as coupling strength, exceeds a critical value at zero temperature QPT1999 . For example, the superradiant phase transition occurs in the Dicke model Dicke when the light-matter coupling strength exceeds a critical value in the thermodynamic limit, i.e., $N\rightarrow\infty$ 1973dickle ; 1973spt ; Sunchangpu ; liyong ; CEPRL ; NLPRL ; ABPRL ; PRL120183603 . The control of the QPT in Dicke model has been discussed ABPRL ; 2023Modulation . Recently, it is found that the QPT can also occur in few-body systems, such as the quantum Rabi model (QRM) 1936Rabi ; 1937Rabi when the ratio of the qubit transition frequency $\omega_{0}$ to the frequency $\omega_{c}$ of the cavity field tends to infinity, i.e., $\eta=\omega_{0}/\omega_{c}\rightarrow\infty$ Hwang2015Rabi . Hereafter, many efforts have been devoted to study the QPT in QRM both theoretically and experimentally trapped ions ; nuclear ; ZhengShiBiao , such as the universal scaling and critical exponents MaoXinLiu2017 , QPT in the quantum Rabi triangle Rabi triangle ; Rabi lattice , dissipative phase transition in the open QRM dispassive QRM ; dar ; sar , QPT in the anisotropic QRM yingzujian ; PRAaqrm and quench dynamics in the anisotropic QRM L. T. ShenPRA . Even the QPT in QRM in the fully quantum limit has also been predicted full quantum limit .
Though the QPT in QRM has been experimentally observed in a single trapped ion trapped ions and a nuclear magnetic resonance quantum simulator nuclear , it is still challenging to manipulate the ground-state QPT due to the limited coupling strength with the classical field (dispersion) limit $\eta\rightarrow\infty$ and the $A^{2}$ terms in cavity quantum electrodynamics (QED) A2 term ; Hwang2015Rabi ; no-go theorm . The $A^{2}$ terms can prevent the occurrence of the QPT in QRM Hwang2015Rabi , which can not be neglected in the ultrastrong-coupling regime in cavity QED system.
In this work, we propose a practical scheme to manipulate the QPT in QRM. Specifically, a periodic modulation was applied to the qubit transition frequency in a standard Rabi model 2017manipulation ; C. Liu in the strong-coupling regime. Then an anisotropic QRM is generated with rotating and counter-rotating coupling strengths allowed to be tuned. Through the modulation, both the effective rotating and counter-rotating (CR) coupling strengths can be varied in a wide range and reach ultrastrong-coupling regime. The ground state is allowed to be in superradiant phase. To investigate the QPT of the anisotropic QRM, we calculated excitation energy and the ground-state phases in both the normal and superradiant phase in the whole coupling parameter space. We find that both the normal and superradiant phase can be reached in the strong-coupling regime via the modulation. We also find that Goldstone modes Hwang2016JC ; Goldstone exist in the superradiant phase. The dependence of excitation energy and ground-state phases on the modulation parameters are analyzed. The feature of the QPT is also investigated by calculating the derivatives of the ground-state energy. Our results show that the QPT in the QRM can be observed and manipulated externally in the strong-coupling regime. These results may inspire other studies on QPT engineering in other few-body systems.
This paper is organized as follows. In Sec. II, we introduce a periodic modulation to the qubit transition frequency in a standard QRM and obtain an effective anisotropic QRM whose rotationg and counter-rotating terms are tunable. In Sec. III, we analyze the dependence of the effective QPT boundary and coupling strengths on modulation parameters. We obtain the excitation energies for five phases and the phase boundary in Sec. IV. In Sec. V, we figure out the phase diagram and verify the phase diagram by analyzing the order parameters. The feature of QPT is also discussed. We also discuss how to manipulate the ground-states phases by varying the modulation parameters in Sec. VI. In Sec. VII, we discuss the $A^{2}$ terms undsr modulation. Finally, we present a brief conclusion and discussions in Sec. VIII.
II Model and Hamiltonian
The system we consider is composed by a two-level system (qubit) coupled to a cavity field, as shown in Fig. 1, which is described by the quantum Rabi model 1936Rabi ; 1937Rabi . The corresponding Hamiltonian reads
$H_{\text{R}}=H_{\mathrm{\text{JC}}}+H_{\mathrm{\text{CR}}},$ (1)
where
$H_{\mathrm{\text{JC}}}=\omega_{c}a^{\dagger}a+\frac{\omega_{0}}{2}\sigma_{z}+g(\sigma_{+}a+a^{\dagger}\sigma_{-})$ (2)
is the usual Jaynes-Cummings (JC) Hamiltonian, and
$H_{\text{CR}}=g(\sigma_{-}a+a^{\dagger}\sigma_{+})$ (3)
describes the CR interactions. Here, $\sigma_{\pm}=(\sigma_{x}\pm i\sigma_{y})$/2 with $\sigma_{x,y,z}$ being the Pauli matrices, and $a\ (a^{\dagger})$ is the annihilation (creation) operator of the cavity field with frequency $\omega_{c}$. We use $\left|e\right\rangle$ and $\left|g\right\rangle$ to label the excited and ground states of the qubit, respectively. The transition frequency of the qubit is $\omega_{0}$ and the coupling strength between the qubit and the cavity field is $g$.
To engineer the system, we introduce a sinusoidal modulation to the transition frequency of the qubit. The modulation Hamiltonian is given by 2017manipulation
$H_{\text{M}}(t)=\frac{1}{2}\xi\nu\cos(\nu t)\sigma_{z},$ (4)
where $\xi$ and $\nu$ are the dimensionless modulation amplitude and frequency, respectively. After including the two-photon $A^{2}$ terms
$H_{A^{2}}=g_{A^{2}}(a+a^{\dagger})^{2}$ (5)
with amplitude $g_{A^{2}}=\chi g^{2}/\omega_{0}$, where $\chi$ is a dimensionless coefficient, the total Hamiltonian can be expressed as
$H(t)=H_{\text{R}}+H_{A^{2}}+H_{\text{M}}(t).$ (6)
To obtain an effective Hamiltonian, we introduce
$H_{0}(t)=\omega_{c}^{\prime}a^{\dagger}a+\frac{1}{2}[\omega_{0}+\xi\nu\cos(\nu t)]\sigma_{z}$ (7)
with $\omega_{c}^{\prime}=\omega_{c}+2g_{A^{2}}$.
Figure 1: (Color online) Schematic of a cavity field coupled to a two-level system (qubit) with a periodic modulation on the qubit transition frequency.
In the first rotating frame defined by
[…]
where $\mathcal{T}$ denotes the time-ordering operator, the Hamiltonian becomes
$\tilde{H}(t)=$ $V_{1}^{\dagger}(t)H(t)V_{1}(t)-iV_{1}^{\dagger}(t)\dot{V}_{1}(t)$
$=$ $\sum_{n=-\infty}^{\infty}gJ_{n}(\xi)\sigma_{+}ae^{i\delta_{n}t}+\sum_{m=-\infty}^{\infty}gJ_{m}(\xi)\sigma_{+}a^{\dagger}e^{i\Delta_{m}t}$
[…]
where we have omitted the constant term in $H_{A^{2}}$. Here, $\delta_{n}=\omega_{0}-\omega_{c}^{\prime}+n\nu$ is the oscillating frequency for the $n$th sideband in the rotating terms, and $\Delta_{m}=\omega_{0}+\omega_{c}^{\prime}+m\nu$ is the oscillating frequency for the $m$th sideband in the CR terms. In the derivation of $\tilde{H}(t)$, we have used the Jacobi-Anger expansion
$e^{i\xi\sin(\nu t)}=\sum_{n=-\infty}^{\infty}J_{n}(\xi)e^{in\nu t}.$ (10)
Here, $J_{n}(\xi)$ is the $n$th Bessel function of the first kind. In Eq. (9), a rotating sideband $n_{0}$ and a CR sideband $m_{0}$ satisfying $\left|gJ_{n_{0}}(\xi)/\delta_{n_{0}}\right|=\max\{\left|gJ_{n}(\xi)/\delta_{n}\right|,n\in Z\}$ and $\left|gJ_{m_{0}}(\xi)/\Delta_{m_{0}}\right|=\max\{\left|gJ_{m}(\xi)/\Delta_{m}\right|,m\in Z\}$ can be selected. Here, “Z” denotes all integers. The oscillating frequencies of all rotating and CR sidebands can be written as $\delta_{n_{0}+s}=\delta_{n_{0}}+s\nu$ and $\Delta_{m_{0}+s}=\Delta_{m_{0}}+s\nu$ with $s$ an integer, respectively. Thus the other sidebands are separated from $n_{0}$ ($m_{0}$) sideband by $s\nu$. Under the conditions
$\nu\gg g>g\left|J_{m}(\xi)\right|,\ g\left|J_{n}(\xi)\right|\ \text{and}\ \nu\gg\left|\Delta_{m_{0}}\right|,\ \left|\delta_{n_{0}}\right|,$ (11)
all the sidebands other than the $n_{0}$($m_{0}$) sideband can be discarded by using RWA. With $g_{A^{2}}\ll g\ll\omega_{0}$ and $g_{A^{2}}\ll 2\omega_{c}^{\prime}$, the $a^{2}$ and $a^{\dagger 2}$ terms with oscillating frequencies ±$\pm 2\omega_{c}^{\prime}$ from the two-photon $A^{2}$ Hamiltonian can also be neglected using RWA. To be focus, the validity of this approximation is discussed in detail in Sec. VII. The Hamiltonian $\tilde{H}(t)$ then can be approximated as
[…]
where we have introduced the effective coupling strengths
$g_{r}(\xi)=gJ_{n_{0}}(\xi),\ g_{cr}(\xi)=gJ_{m_{0}}(\xi).$ (13)
The Hamiltonian describes an effective standard Rabi model with $g_{r}=g_{cr}$ or an effective anisotropic QRM with $g_{r}\neq g_{cr}$ QTxie in interaction picture. Note that, we can tune the effective rotating and CR coupling strengths $g_{r}(\xi)$ and $g_{cr}(\xi)$ via the dimensionless modulation amplitude $\xi$. To make Hamiltonian (12) time independent, we introduce the effective frequency for the qubit
$\tilde{\omega}_{0}$ $=\frac{\Delta_{m_{0}}+\delta_{n_{0}}}{2},$
and the cavity field
$\tilde{\omega}_{c}=\frac{\Delta_{m_{0}}-\delta_{n_{0}}}{2}.$
After transferring to the second rotating frame by the unitary transformation
$V_{2}(t)=e^{-i(\frac{\tilde{\omega}_{0}}{2}\sigma_{z}+\tilde{\omega}_{c}a^{\dagger}a)t},$ (14)
the Hamiltonian (12) becomes
${H}_{eff}$ $=V_{2}(t)\tilde{H}_{1}(t)V_{2}^{\dagger}(t)+i\dot{V}_{2}(t)V_{2}^{\dagger}(t)$
[…]
Here, we have introduced the anisotropic parameter $\epsilon(\xi)=g_{cr}(\xi)/g_{r}(\xi)$. The effective coupling strength $g_{r}(\xi)$, the effective frequencies $\tilde{\omega}_{0}$ and $\tilde{\omega}_{c}$, and anisotropic parameter $\epsilon(\xi)$ can be regulated by the modulation parameters $\nu$ and $\xi$. The effective Hamiltonian (15) describe an anisotropic QRM with $\epsilon(\xi)\neq 1$ and a standard QRM with $\epsilon(\xi)=1$. Hereafter, we write $\epsilon(\xi)$ as $\epsilon$ for short.
III Ultrastrong coupling
The effective standard quantum Rabi model ($\epsilon=1$) in the limit $\tilde{\eta}=\tilde{\omega}_{0}/\tilde{\omega}_{c}\rightarrow\infty$ can exhibit quantum phase transition from the normal phase to the superradiant phase. This phase transition occurs at $g_{C}=\sqrt{\tilde{\omega}_{0}\tilde{\omega}_{c}}/2$ Ashhab2013 above which $Z_{2}$ parity symmetry is broken and the bosonic mode is macroscopically occupied Hwang2015Rabi . Here, the critical coupling strength $g_{C}$ depends on the modulation frequency $\nu$. Since the occurrence of the QPT in the quantum Rabi model requires the limit $\tilde{\eta}\rightarrow\infty$, to see clearly the dependence of $\tilde{\eta}$ on the modulation parameters, we plot $\tilde{\eta}$ versus $\nu$ for different ratios $\eta=\omega_{0}/\omega_{c}^{\prime}$ in Fig. 2(a). $\tilde{\eta}$ is independent of $\xi$. As shown in Fig. 2(a), the curve of the effective ratio $\tilde{\eta}$ is composed by series of V-type valleys until $\nu$ increases up to $\nu/\omega_{0}=2(\omega_{0}+\omega_{c}^{\prime})$. The V-shaped valleys share the same $m_{0}$ and the valley for $m_{0}$ starts at $\nu=2(\omega_{0}+\omega_{c}^{\prime})/(2\left|m_{0}\right|+1)$. When $\nu>2(\omega_{0}+\omega_{c}^{\prime})$, $m_{0}=0$ and $\tilde{\eta}=\eta$. Thus, in the regime $\nu<2(\omega_{0}+\omega_{c}^{\prime})$, we can manipulate $\tilde{\eta}$ via choosing appropriate $\nu$ while in the regime $\nu>2(\omega_{0}+\omega_{c}^{\prime})$, this $\tilde{\eta}$ can not be varied via changing $\nu$. To see clearly the dependence of the effective coupling strengths $g_{r}(\xi)$ and $g_{cr}(\xi)$ on the modulation amplitude $\xi$, we plot $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ versus the modulation amplitude $\xi$ for $\eta=1$ and $10^{2}$ at $\nu/\omega_{0}=0.49$ and $0.68$ in Figs. 2(b) and 2(c), respectively. The effective coupling strength $g_{r}/g_{C}\ (g_{cr}/g_{C})$ depends on the modulation amplitude $\xi$ in the form of the Bessel function. When $\nu/\omega_{0}=0.49$ for $\eta=1$ $(m_{0}=-4,\ n_{0}=-1)$, $g_{r}/g_{C}$ oscillates between -2.41656 and 6 with reducing amplitude, and $g_{cr}/g_{C}$ oscillates between -1.64629 and 2.39791, as shown in Fig. 2(b). When $\nu/\omega_{0}=0.68$ for $\eta=10^{2}$ $(m_{0}=n_{0}=-1)$, $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ show the same oscillating feature between -1.23432 and 0.734245, as shown in Fig. 2(c). Note that, $g_{r}=g_{cr}$ in Fig. 2(c) under the chosen parameters. The results show that one can tune the effective coupling strengths $g_{r}$ and $g_{cr}$ in a broad range.
Figure 2: (Color online) (a) The effective ratio $\tilde{\eta}$ versus the modulation frequency $\nu/\omega_{0}$ for $\eta=10^{2}\text{ and }10^{3}.$ (b-c) The couplings $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ versus the modulation amplitude $\xi$. Here, $\nu/\omega_{0}=0.49$ and $\eta=1$ in panel (b) and $\nu/\omega_{0}=0.68$ and $\eta=10^{2}$ in panel (c), other parameter is $g/\omega_{0}=0.06.$
IV Excitation spectra
To figure out the ground-state property of the system in the whole coupling-parameter space, we calculate the excitation energy in the limit $\tilde{\eta}\rightarrow\infty$ in this section.
IV.1 Normal phase
To diagonalize the effective Hamiltonian (15), we introduce the canonical coordinate and momentum operators $x=(a+a^{\dagger})/\sqrt{2}$ and $p=i(a^{\dagger}-a)/\sqrt{2}$. In terms of $x$ and $p$, the Hamiltonian (15) reads
$H_{\text{AR}}=\tilde{H}_{\text{0}}+\tilde{H}_{\text{I}},$ (16)
where
$\tilde{H}_{0}$ $=\frac{\tilde{\omega}_{c}}{2}(x^{2}+p^{2})+\frac{\tilde{\omega}_{0}}{2}\sigma_{z},$ (17a)
$\tilde{H}_{\text{I}}$ $=\frac{g_{r}}{\sqrt{2}}[(1+\epsilon)\sigma_{x}x-(1-\epsilon)\sigma_{y}p].$ (17b)
Then, in the limit $\tilde{\eta}\rightarrow\infty$, we make an unitary transformation $U_{\text{N}}^{\dagger}{H}_{\text{AR}}U_{\text{N}}$ on Hamiltonian (16) with
$U_{\text{N}}=\exp\{-ig_{r}[(1-\epsilon)\sigma_{x}p+(1+\epsilon)\sigma_{y}x]/{\sqrt{2}\tilde{\omega}_{0}}\}.$ (18)
The transformed Hamiltonian becomes free of coupling between qubit subspaces $\mathcal{H}_{g}$ and $\mathcal{H}_{e}$, which are spanned by $\left|g\right\rangle$ and $\left|e\right\rangle$, respectively. After a projection into $\mathcal{H}_{g}$, i.e., $\tilde{H}_{\text{N}}=\langle g|U_{\text{N}}^{\dagger}H_{\text{AR}}U_{\text{N}}|g\rangle$, we obtain an effective low-energy Hamiltonian as
$\tilde{H}_{\text{N}}=\frac{{\tilde{\omega}_{c}}}{2}[(1-\zeta^{2})x^{2}+(1-\zeta^{\prime^{2}})p^{2}],$ (19)
where a constant term is neglected. Here,
Figure 3: (Color online) The phase transition boundary $\tilde{g}_{C}/\omega_{0}$ for the anisotropic QRM versus $\xi$ for $\nu/\omega_{0}=0.402$ and $0.68$ with $\eta=10^{2}$.
$\zeta=\frac{\lambda(1+\epsilon)}{2}\ \text{and}\ \zeta^{\prime}=\frac{\lambda(1-\epsilon)}{2}$ (20)
with the dimensionless rotating-coupling strength $\lambda\equiv{g_{r}/g_{C}}$. We note that there is no interaction term in Eq. (19). The excitation energy $\omega_{\text{N}}$ can be directly obtained as
$\omega_{\text{N}}=\tilde{\omega}_{c}\sqrt{(1-\zeta^{2})(1-\zeta^{\prime 2})}.$ (21)
$\tilde{H}_{\text{N}}$ implies divergent values of $\langle x\rangle$ ($\langle p\rangle$) due to the disappearance of $x^{2}$ ($p^{2}$) term when $\epsilon>0$ ($\epsilon<0$) in the limit $\tilde{\eta}\rightarrow\infty$. By this fact, the normal phase is restricted in the region
$g_{r}<\left|\tilde{g}_{C}\right|$ (22)
with
$\tilde{g}_{C}=2g_{C}/(1+\left|\epsilon\right|).$ (23)
In the normal phase, the value of $\langle x\rangle\ (\langle p\rangle)$ is zero. $\tilde{g}_{C}$ determines the phase transition boundary for the anisotropic QRM. Note that, $g_{C}$ is the critical coupling strength for the standard QRM. When $g_{r}<\left|\tilde{g}_{C}\right|$, the ground state is in the normal phase with the cavity field in vacuum and qubit in $\left|g\right\rangle$. While outside the normal phase, namely, $g_{r}>\left|\tilde{g}_{C}\right|$, the system is in the superradiant phase with both the qubit and the cavity field excited.
In Fig. 3, we plot $\tilde{g}_{C}/\omega_{0}$ versus $\xi$ for $\nu/\omega_{0}=0.402\ \text{and}\ 0.68$. For $\nu/\omega_{0}=0.402$, $\tilde{g}_{C}/\omega_{0}$ oscillates between 0 and 0.0309031. The zero points of $\tilde{g}_{C}/\omega_{0}$ locate at those zeros of $J_{n_{0}}(\xi)$. The first zero point of $\tilde{g}_{C}/\omega_{0}$ locates at $\xi=5.13562$. For a larger $\nu$, such as $\nu/\omega_{0}=0.68$, $\tilde{g}_{C}/\omega_{0}$ becomes independent of $\xi$ due to $m_{0}=n_{0}$. Thus, we can tune the QPT boundary via $\nu$ and $\xi$.
IV.2 Superradiant phase
Now, we study the excitation energy in the superradiant phase, namely, $g_{r}>\left|\tilde{g}_{C}\right|$. In the superradiant phase, there are four different superradiant phases. When $\epsilon>0$, the superradiant phase can be defined as superradiant X (SX) phase since the $\langle x\rangle$ becomes divergent. When $\epsilon<0$, the system is in superradiant P (SP) phase since the $\langle p\rangle$ is divergent. For $g_{cr}>2\left|g_{C}\right|$ with $g_{r}=0$, the ground state of system is in superradiant XP a (SXPa) phase while for $g_{r}>2\left|g_{C}\right|$ with $g_{cr}=0$, the system is in superradiant XP b (SXPb) phase. When $g_{cr}=0$, only rotating terms exist in the Hamiltonian, which reduces to a Jaynes-Cummings model. In the following, we investigate the excitation energy in the limit $\tilde{\eta}\rightarrow\infty$ for SX, SP, SXPa and SXPb phase, respectively. To effectively describe the superradiant phase with macroscopic excitation, we displace the cavity field $a$ by displacing operator $D_{k}(\alpha)$ (k=SX, SP, SXPa, SXPb) for k phase. After the displacement, the Hamiltonian $H_{\text{AR}}$ becomes
$\tilde{H}_{AR}=D_{k}^{\dagger}(\alpha)H_{\text{AR}}D_{k}(\alpha).$ (24)
Here, the detailed form of $D_{k}(\alpha)$ is given in the Appendix A. To diagonalize the qubit part in Hamiltonian $\tilde{H}_{AR}$, we introduce a new set of Pauli operators $\tau_{x,y,z}^{k}$ (k=SX, SP, SXPa, SXPb) for each k phase, respectively. To be concise, the form of $\tau_{x,y,z}^{k}$ for k phase is also given in the Appendix A. In terms of the new Pauli operators $\tau_{x,y,z}^{k}$ , the displaced Hamiltonian $\tilde{H}_{\text{AR}}$ becomes
$\tilde{H}_{\text{AR}}(\alpha_{\text{SX}})=$ $H_{c}+\frac{\Omega_{\text{SX}}}{2}\tau_{z}^{{}_{\text{SX}}}-\chi_{2}(\zeta^{\prime})\tau_{y}^{{}_{\text{SX}}}p+\chi_{1}(\zeta)\tau_{x}^{{}_{\text{SX}}}x$ (25)
$+\frac{\tilde{\omega}_{c}}{2}\alpha_{\text{SX}}^{2},$
$\tilde{H}_{\text{AR}}(\alpha_{\text{SP}})=$ $H_{c}+\frac{\Omega_{\text{SP}}}{2}\tau_{z}^{{}_{\text{SP}}}-\chi_{1}(\zeta^{\prime})\tau_{x}^{{}_{\text{SP}}}p-\chi_{2}(\zeta)\tau_{y}^{{}_{\text{SP}}}x$
$+\frac{\tilde{\omega}_{c}}{2}\alpha_{\text{SP}}^{2},$
$\tilde{H}_{\text{AR}}(\alpha_{\text{SXPa}})=$ $\frac{\Omega_{\text{SXPa}}}{2}\tau_{z}^{{}_{\text{SXPa}}}+\chi_{3}(\frac{4}{\mu}\tau_{x}^{{}_{\text{SXPa}}}x+\mu\tau_{y}^{{}_{\text{SXPa}}}p)$
$+H_{c}+\tilde{\omega}_{c}\alpha_{\text{SXPa}}^{2},$
$\tilde{H}_{\text{AR}}(\alpha_{\text{SXPb}})=$ $\frac{\Omega_{\text{SXPb}}}{2}\tau_{z}^{{}_{\text{SXPb}}}+\chi_{3}(\frac{4}{\lambda}\tau_{x}^{{}_{\text{SXPb}}}x-\lambda\tau_{y}^{{}_{\text{SXPb}}}p)$
$+H_{c}+\tilde{\omega}_{c}\alpha_{\text{SXPb}}^{2}.$
Here, the free cavity field Hamiltonian $H_{c}$ is
$H_{c}=\frac{\tilde{\omega}_{c}}{2}(x^{2}+p^{2})-\frac{\tilde{\omega}_{c}}{2},$ (26)
and the rescaled qubit frequency $\Omega_{k}$ for k phase is
$\begin{array}[]{r@{}l}\Omega_{\text{SX}}&{}=\tilde{\omega}_{0}\zeta^{2},\\ \Omega_{\text{SP}}&{}=\tilde{\omega}_{0}\zeta^{\prime 2},\\ \Omega_{\text{SXPa}}&{}=g_{cr}^{2}/\tilde{\omega}_{c},\\ \Omega_{\text{SXPb}}&{}=g_{r}^{2}/\tilde{\omega}_{c}.\end{array}$ (27)
The displacement parameter $\alpha$ is determined by the vanishment of the non-diagonal terms during deriving Eq. (25), which is obtained as
$\alpha_{\text{SX}}$ =±$=\pm\sqrt{\frac{\tilde{\eta}}{2}(\zeta^{2}-\frac{1}{\zeta^{2}})},$ (28)
$\alpha_{\text{SP}}$ =±$=\pm\sqrt{\frac{\tilde{\eta}}{2}(\zeta^{\prime 2}-\frac{1}{\zeta^{\prime 2}})},$
$\alpha_{\text{SXPa}}$ =±$=\pm\sqrt{\tilde{\eta}(\frac{\mu^{2}}{16}-\frac{1}{\mu^{2}})},$
$\alpha_{\text{SXPb}}$ =±$=\pm\sqrt{\tilde{\eta}(\frac{\lambda^{2}}{16}-\frac{1}{\lambda^{2}})},$
for k (k=SX, SP, SXPa, SXPb) phase. To distinguish $\alpha$ in each k phase, here we have added a subscript “k” in $\alpha$, i.e., we label $\alpha$ as $\alpha_{k}$. Here, the dimensionless CR coupling strength $\mu$ is defined as $\mu=g_{cr}/g_{C}$. The parameter $\chi_{i}$ (i=1,2,3) in Eq. (25) is defined as
$\chi_{1}(\zeta)=\frac{\sqrt{2}g_{C}}{\zeta},\ \ \chi_{2}(\zeta)=\sqrt{2}g_{C}\zeta,\ \text{and}\ \chi_{3}=\frac{g_{C}}{\sqrt{2}}.$ (29)
We find that the form of the transformed Hamiltonian $\tilde{H}_{\text{AR}}(\alpha_{k})$ in Eq. (25) is similar to Hamiltonian (16) with rescaled qubit frequency $\Omega_{k}$.
[…]
Thus, we can make similar unitary transformation $U_{k}$ as $U_{\text{N}}$ in the normal phase on $\tilde{H}_{\text{AR}}(\alpha_{k})$ to eliminate the non-diagonal terms, namely,
$\tilde{H}_{k}=U_{k}^{\dagger}\tilde{H}_{\text{AR}}(\alpha_{k})U_{k},$ (30)
(k=SX, SP, SXPa, SXPb). Here, the detailed form of $U_{k}$ is given in the Appendix A. After performing the similar procedure used to derive $\tilde{H}_{\text{N}}$, we obtain the effective Hamiltonian $\tilde{H}_{k}$ in k phase as following:
$\tilde{H}_{\text{SX}}$ $=\frac{\tilde{\omega}_{c}}{2}(1-\frac{1}{\zeta^{4}})x^{2}+\frac{\tilde{\omega}_{c}}{2}(1-\frac{\zeta^{\prime 2}}{\zeta^{2}})p^{2},$ (31)
$\tilde{H}_{\text{SP}}$ $=\frac{\tilde{\omega}_{c}}{2}(1-\frac{\zeta^{2}}{\zeta^{\prime 2}})x^{2}+\frac{\tilde{\omega}_{c}}{2}(1-\frac{1}{\zeta^{\prime 4}})p^{2},$
$\tilde{H}_{\text{SXPa}}$ $=\frac{\tilde{\omega}_{c}}{4}(1-\frac{16}{\mu^{4}})x^{2}-\frac{\tilde{\omega}_{c}}{2}(1+\frac{4}{\mu^{2}})-\tilde{\omega}_{0}(\frac{\mu^{2}}{16}+\frac{1}{\mu^{2}}),$
$\tilde{H}_{\text{SXPb}}$ $=\frac{\tilde{\omega}_{c}}{4}(1-\frac{16}{\lambda^{4}})x^{2}-\frac{\tilde{\omega}_{c}}{2}(1-\frac{4}{\lambda^{2}})-\tilde{\omega}_{0}(\frac{\lambda^{2}}{16}+\frac{1}{\lambda^{2}}).$
Note that Hamiltonian $\tilde{H}_{k}$ in Eq. (31) describes a free bosonic mode. Interestingly, the $p^{2}$ term disappears in $\tilde{H}_{\text{SXPa}}$ and $\tilde{H}_{\text{SXPb}}$. Now, the excitation energy $\omega_{k}$ for each k phase can be obtained as
$\begin{array}[]{r@{}l}\omega_{\text{SX}}&{}=\tilde{\omega}_{c}\sqrt{(1-\frac{1}{\zeta^{4}})(1-\frac{\zeta^{\prime 2}}{\zeta^{2}})},\\ \omega_{\text{SP}}&{}=\tilde{\omega}_{c}\sqrt{(1-\frac{1}{\zeta^{\prime 4}})(1-\frac{\zeta^{2}}{\zeta^{\prime 2}})},\\ \omega_{\text{SXPa}}&{}=0,\\ \omega_{\text{SXPb}}&{}=0.\end{array}$ (32)
We find that the excitation energy closes both in SXPa and SXPb phases. This result denotes the existence of the Goldstone modes Hwang2016JC ; Goldstone . Furthermore, $\omega_{\text{{SP}}}$ has the same form as $\omega_{\text{{SX}}}$ if making the replacement $\zeta\leftrightarrow$$\zeta^{\prime}$ in $\omega_{\text{{SX}}}$.
[…]
V Phase diagram
In this section, we investigate the phase diagram based on the above analysis and order parameters. Then we discuss the feature of QPTs by studying the discontinuity of the derivatives of the ground-state energy $E_{G}$.
Figure 5: (Color online) $\langle x\rangle$ in panel (a) and $\langle p\rangle$ in panel (b) versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ for $\eta=10^{2}$.
To study the phase diagram, we introduce the average values $\langle x\rangle$ and $\langle p\rangle$ as the order parameters. The average values $\langle x\rangle$ and $\langle p\rangle$, and the corresponding variances $\Delta x^{2}$ and $\Delta p^{2}$ for p phase are obtained from the ground state of Hamiltonian $\tilde{H}_{p}$ (p=N, SX, SP, SXPa, SXPb). We show the results in table 1. The expressions of $s$, $v$ and $w$ in table 1 are defined as
[…]
$v(x_{1},x_{2})$ $=\frac{1}{2}\sqrt{\frac{1-x_{2}^{2}}{1-x_{1}^{2}}},$ (33b)
$w(x_{1},x_{2})$ $=\frac{\sqrt{x_{1}}}{1+x_{1}}\frac{x_{2}^{2}}{\sqrt{x_{2}^{4}-1}}.$ (33c)
We find that, in the limit $\tilde{\eta}\rightarrow\infty$, $\langle x\rangle\rightarrow\infty$ and $\langle p\rangle=0$ in SX phase while $\langle x\rangle=0$ and $\langle p\rangle\rightarrow\infty$ in the SP phase. That is why they are defined as SX phase and SP phase, respectively. We also note that, the variances $\Delta x^{2}$ and $\Delta p^{2}$ are finite for N, SX and SP phases, but $\Delta x^{2}=0$ and $\Delta p^{2}=\infty$ for both SXPa and SXPb phases.
Table 1: $\langle x\rangle$, $\langle p\rangle$, $\langle x^{2}\rangle$ and $\langle p^{2}\rangle$ for all phases.
N SX SP SXPa SXPb
$\langle x\rangle$ 0 $s(\zeta)$ 0 $s(\frac{\mu}{2})$ $s(\frac{\lambda}{2})$
$\langle p\rangle$ 0 0 $s(\zeta^{\prime})$ 0 0
$\Delta x^{2}$ $v(\zeta,\zeta^{\prime})$ $w(\epsilon,\zeta)$ $4^{-1}w^{-1}(-\epsilon,\zeta^{\prime})$ 0 0
$\Delta p^{2}$ $v(\zeta^{\prime},\zeta)$ $4^{-1}w^{-1}(\epsilon,\zeta)$ $w(-\epsilon,\zeta^{\prime})$infty $\infty$
To see clearly the dependence of $\langle x\rangle$ and $\langle p\rangle$ on the coupling strengths $g_{r}$ and $g_{cr}$, we plot $\langle x\rangle$ and $\langle p\rangle$ versus the coupling strengths $g_{r}$ and $g_{cr}$ in Fig. 5. In the region $g_{r}<\left|\tilde{g}_{C}\right|$, both $\langle x\rangle$ and $\langle p\rangle$ are 0, which correspond to the normal phase. Outside the normal phase region, i.e., $g_{r}>\left|\tilde{g}_{C}\right|$, when $\epsilon>0$, the ground state is in SX phase with $\langle x\rangle>0$; when $\epsilon<0$, the ground state is in SP phase with $\langle p\rangle>0$. One can distinguish the phases by $\langle x\rangle$ and $\langle p\rangle$.
Figure 6: (a) (Color online) (a-b) The first derivatives $\partial E_{G}/\partial\lambda$ and $\partial E_{G}/\partial\mu$ versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$. (b-c) The second derivatives $\partial^{2}E_{G}/\partial\lambda^{2}$ and $\partial^{2}E_{G}/\partial\mu^{2}$ versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$. Other parameters are $\omega_{0}=1,\eta=10^{2}$.
To characterize the type of the QPTs, we calculate the derivatives of ground-state energy $E_{G}$ over the coupling strengths $g_{r}$ and $g_{cr}$. To start with, we first calculate the ground-state energy $E_{G}$ in all phases. We perform the squeezing transformation
$S(r_{p})=\exp[\frac{r_{p}}{2}(a^{2}-a^{\dagger 2})]$ (34)
on $\tilde{H}_{p}$ (p=N, SX, SP, SXPa and SXPb). The detailed form of $r_{p}$ is given in the Appendix B. Then the resulting Hamiltonian reads
$\tilde{H}(r_{p})=$ $S^{\dagger}(r_{p})\tilde{H}_{p}S(r_{p})$
$=$ $\omega_{p}a^{\dagger}a+E_{G}^{p},$ (35)
where $\omega_{p}$ is the excitation energy given by Eqs. (21) and (32), and $E_{G}^{p}$ is the ground-state energy in p phase, which is obtained as
$E_{G}^{\text{N}}=$ $\frac{\omega_{\text{N}}}{2}+\frac{\tilde{\omega}_{c}}{2}\zeta\zeta^{\prime}-\frac{\tilde{\omega}_{0}}{2}-\frac{\tilde{\omega}_{c}}{2},$ (36)
$E_{G}^{\text{SX}}=$ $\frac{\omega_{\text{SX}}}{2}+\frac{\tilde{\omega}_{c}}{2}\frac{\zeta^{\prime}}{\zeta^{3}}-\frac{\tilde{\omega}_{0}}{4}(\zeta^{2}+\zeta^{-2})-\frac{\tilde{\omega}_{c}}{2},$
$E_{G}^{\text{SP}}=$ $\frac{\omega_{\text{SP}}}{2}+\frac{\tilde{\omega}_{c}}{2}\frac{\zeta}{\zeta^{\prime 3}}-\frac{\tilde{\omega}_{0}}{4}(\zeta^{\prime 2}+\zeta^{\prime-2})-\frac{\tilde{\omega}_{c}}{2},$
$E_{G}^{\text{SXPa}}=$ $\-\frac{\tilde{\omega}_{c}}{2}(1+\frac{4}{\mu^{2}})-\tilde{\omega}_{0}(\frac{\mu^{2}}{16}+\frac{1}{\mu^{2}}),$
$E_{G}^{\text{SXPb}}=$ $\frac{\tilde{\omega}_{c}}{2}(1-\frac{4}{\lambda^{2}})-\tilde{\omega}_{0}(\frac{\lambda^{2}}{16}+\frac{1}{\lambda^{2}}).$
We plot the dependence of the first and second derivatives of $E_{G}$ on $\lambda$ and $\mu$ versus $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ in Fig. 6. Note that, $\lambda$ and $\mu$ are dimensionless rotating and CR coupling strengths, respectively. As shown in Fig. 6(a) and 6(b), it is found that $\partial E_{G}/\partial\lambda$ ($\partial E_{G}/\partial\mu$) is discontinuous at $g_{r}=0\ (g_{cr}=0)\ \text{for}\ \left|g_{cr}/g_{C}\right|>2$ ($\left|g_{r}/g_{C}\right|>2$). This results denote the QPT from SX to SP is of first order. As shown in Figs. 6(c) and 6(d), we observe that $\partial^{2}E_{G}/\partial\lambda^{2}$ ($\partial^{2}E_{G}/\partial\mu^{2}$) is discontinuous at the boundaries $g_{r}=\left|\tilde{g}_{C}\right|$ and $g_{r}=0$ ($g_{cr}=0$) for $\left|g_{cr}/g_{C}\right|>2$ ($\left|g_{r}/g_{C}\right|>2$). Thus, the transition from the normal phase to superradiant phases is of second order. We also use the dashed and solid lines to indicate the first and second order QPT in Fig. 5, respectively.
VI Manipulation of Quantum phases
In this section, we discuss how to manipulate the ground state by varying the modulation parameters. We plot the excitation energy $\omega$ as a function of $\xi$ at $\nu/\omega_{0}=0.68$ in Fig. 7. The subscript “p” that labels the specific phase in $\omega_{p}$ is omitted.
Figure 7: (Color online) The excitation energy $\omega$ versus the dimensionless modulation amplitude $\xi$. Parameters are $\eta=10^{2},\nu/\omega_{0}=0.68$ and $g/\omega_{0}=0.06$.
The solid curve indicates that the system can enter different phases by adjusting the dimensionless modulation amplitude $\xi$. The QPT occurs at $\xi=1.10172$ and 2.59789, where $g_{r}/g_{C}=-1$, $g_{cr}/g_{C}=-1$. The ground state is in the normal phase when $1.101720<\xi<$ 1.10172 and $32.59789<\xi<$ 3, and the SX phase appears when $2.597891.10172<\xi<$ 2.59789. The system experiences the QPTs from the normal phase to SX phase at $\xi=1.10172$ and from SX phase to the normal phase at $\xi=2.59789$.
To see the system how to cross the phase boundary under the chosen parameter in Fig. 7, we plot the parameterized curves of [$g_{r}(\xi)/g_{C}$, $g_{cr}(\xi)/g_{C}$] versus the modulation amplitude $\xi$ in Fig. 8(a). The parameterized curve evolves from the normal phase to SX phase then back to the normal phase as $\xi$ increases from 0 to 3. The system can be described by a standard QRM. In Fig. 8(b), we plot $\langle x\rangle$ versus the modulation amplitude $\xi$ for the same parameters as Fig. 8(a). Note that, $\langle p\rangle=0$ under the parameters in Fig. 8(a). As shown in Fig. 8(b), although the coupling strength is only in the strong-coupling regime with $g/\omega_{0}=0.06$, yet by the modulation, the QRM can reach the superradiant phase when $2.597891.10172<\xi<$ 2.59789. The normal phase appears when $1.101720<\xi<$ 1.10172 and $32.59789<\xi<$ 3. The results shown in Fig. 8 are consistent with those in Fig. 7. Thus, we can manipulate quantum phases on the QRM via the appropriate modulation with strong coupling.
Figure 8: (Color online) (a) The parameterized curve of the couplings $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ with the modulation amplitude $\xi$ increasing from 0 to 3 as indicated by the arrows. (b) $\langle x\rangle$ versus the modulation amplitude $\xi$. The parameters are the same as Fig. 7.
We mention that, the anisotropic QRM can be realized with our scheme, as shown in Fig. 2(b). However, to reach the classical field limit $\tilde{\eta}\rightarrow\infty$, only an effective standard QRM can be realized. This is because in the limit $\tilde{\eta}\rightarrow\infty$ and the condition (11), we obtain $m_{0}=n_{0}$, thus $g_{r}=g_{cr}$. This is an effective standard QRM. Despite this, we still provide a possible scheme to observe the superradiant phase of few-body quantum systems in the strong-coupling regime.
In this paper, we consider negligible cavity dissipation with $\kappa/\omega_{0}<10^{-4},\ \kappa/\omega_{c}^{\prime}<10^{-2}$. When the cavity dissipation is non-negligible, it will produce a shift to the critical coupling strength, which becomes $g_{C}^{\text{diss}}=\frac{1}{2}\sqrt{\frac{\tilde{\omega}_{0}}{\tilde{\omega}_{c}}(\tilde{\omega}_{c}^{2}+\kappa^{2})}$. For a dissipation rate $\kappa=0.2\omega_{c}^{\prime}=0.002\omega_{0}$ and $\nu/\omega_{0}=0.68$, $\tilde{\omega}_{c}=0.01\omega_{0}$ and $g_{C}^{\text{diss}}\approx 1.02g_{C}$. The effective coupling strengths $g_{r}$ and $g_{cr}$ can be larger than $g_{C}^{\text{diss}}$, as shown in Fig. 2(c). Hence, it is still effective for our approach to manipulate the steady-state superradiant phase transition of the dissipative QRM.
VII two-photon $A^{2}$ terms
In this section, we discuss the validity of omitting the two-photon $A^{2}$ terms. The two-photon $A^{2}$ terms given by Hamiltonian (5) can be written as
$H_{A^{2}}=2g_{A^{2}}a^{\dagger}a+g_{A^{2}}(a^{\dagger 2}+a^{2}).$ (37)
The first term in $H_{A^{2}}$ has been absorbed into the Hamiltonian of the standard QRM by replacing the bosonic mode frequency $\omega_{c}$ with the modified frequency $\omega_{c}^{\prime}=\omega_{c}+2g_{A^{2}}$. The second term can be safely neglected with $g\ll\omega_{0},\ g_{A^{2}}\ll 2\omega_{c}^{\prime}$. To demonstrate this, in Fig. 9, we plot the ratio $g_{A^{2}}/2\omega_{c}^{\prime}$ versus the coupling strength $g$ and the dimensionless coefficient $\chi$. For the dispersion case $\eta=10^{2}$, as shown in Fig. 9(a), to satisfy $g_{A^{2}}/2\omega_{c}^{\prime}=0.01$, as indicated by the dashed curve, $g$ and $\chi$ need to be ($g/\omega_{0}\sim 0.06,\ \chi\sim 0.0579$) or ($g/\omega_{0}\sim 0.0102,\ \chi\sim 2$). For the parameters we used in our discussions with $\eta=10^{2}$, $g/\omega_{0}=0.06$, and $\chi\leq 0.05$, this ratio is smaller than $1\%$, namely, $g_{A^{2}}/2\omega_{c}^{\prime}<0.01.$ Hence, we can neglect these terms in the dispersion limit. For the resonance case $\eta=1$, as shown in Fig. 9(b), to satisfy $g_{A^{2}}/2\omega_{c}^{\prime}=0.01$, we should have ($g/\omega_{0}\sim 0.06,\ \chi\sim 5.787$). Therefore, it is also reasonable to omit the $A^{2}$ terms for $g/\omega_{0}=0.06$ and $\chi<5$ for resonance case $\eta=1$. Thus we can omit the $A^{2}$ terms from the following discussions.
Figure 9: (Color online) The ratio $g_{A^{2}}/2\omega_{c}^{\prime}$ versus the coupling strength $g/\omega_{0}$ and the dimensionless factor $\chi$. Parameters are (a) $\eta=10^{2}$ and (b) $\eta=1$.
VIII conclusion and discussions
By applying a qubit frequency modulation to a standard QRM, we generate an anisotropic QRM whose effective couplings for rotating and counter-rotating terms can be adjusted over a wide range. We calculate the ground-state and excitation energy of this model in terms of the modulation parameters. The phase diagram of the anisotropic QRM in the whole coupling parameter space and the type of the QPT are also obtained. We show that the superradiant phase of the QRM can be observed in the strong-coupling regime with the frequency modulation. Furthermore, the two-photon $A^{2}$ terms, which could prevent the observation of the QPT in QRM Hwang2015Rabi , can be appropriately neglected in our scheme. We provide an experimentally accessible scheme to control the QPT in QRM. Our scheme is within current experimental technology trapped ions ; nuclear ; ZhengShiBiao ; 24.2 . This modulation scheme may also be applicable in other few-body quantum systems. Our work may pave the way for experimentally observing superradiant phase in other few-body quantum systems.
Acknowledgements.
J.-F.H. is supported in part by the National Natural Science Foundation of China (Grant No. 12075083).
Appendix A Operators for superradiant phase
The form of displacing operator $D_{k}(\alpha)$ (k=SX, SP, SXPa, SXPb) for each k phase reads
$\begin{array}[]{r@{}l}D_{\text{SX}}(\alpha)&{}=e{}^{-\alpha\frac{\partial}{\partial x}},\\ D_{\text{SP}}(\alpha)&{}=e^{-\alpha\frac{\partial}{\partial p}},\\ D_{\text{SXPa}}(\alpha)&{}=e^{-\sqrt{2}\alpha\frac{\partial}{\partial x}},\\ D_{\text{SXPb}}(\alpha)&{}=e^{-\sqrt{2}\alpha\frac{\partial}{\partial x}}.\end{array}$ (38)
The new Pauli operators we introduced are
$\begin{array}[]{r@{}l}\tau_{x}^{\text{SX}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SX}}}\sigma_{x}-\frac{1}{\Omega_{\text{SX}}}\sqrt{2\tilde{\omega}_{0}\tilde{\omega}_{c}}\zeta\left|\alpha_{\text{SX}}\right|\sigma_{z},\\ \tau_{y}^{\text{SX}}&{}=\sigma_{y},\\ \tau_{z}^{\text{SX}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SX}}}\sigma_{z}+\frac{1}{\Omega_{\text{SX}}}\sqrt{2\tilde{\omega}_{0}\tilde{\omega}_{c}}\zeta\left|\alpha_{\text{SX}}\right|\sigma_{x}\end{array}$ (39)
for SX phase,
$\begin{array}[]{r@{}l}\tau_{x}^{\text{SP}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SP}}}\sigma_{y}+\frac{1}{\Omega_{\text{SP}}}\sqrt{2\tilde{\omega}_{0}\tilde{\omega}_{c}}\zeta^{\prime}\left|\alpha_{\text{SP}}\right|\sigma_{z},\\ \tau_{y}^{\text{SP}}&{}=-\sigma_{x},\\ \tau_{z}^{\text{SP}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SP}}}\sigma_{z}-\frac{1}{\Omega_{\text{SP}}}\sqrt{2\tilde{\omega}_{0}\tilde{\omega}_{c}}\zeta^{\prime}\left|\alpha_{\text{SP}}\right|\sigma_{y}\end{array}$ (40)
for SP phase,
$\begin{array}[]{r@{}l}\tau_{x}^{\text{SXPa}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPa}}}\sigma_{x}-\frac{1}{\Omega_{\text{SXPa}}}2g_{c}\left|\alpha_{\text{SXPa}}\right|\sigma_{z},\\ \tau_{y}^{\text{SXPa}}&{}=\sigma_{y},\\ \tau_{z}^{\text{SXPa}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPa}}}\sigma_{z}+\frac{1}{\Omega_{\text{SXPa}}}2g_{c}\left|\alpha_{\text{SXPa}}\right|\sigma_{x}\end{array}$ (41)
for SXPa phase and
$\begin{array}[]{r@{}l}\tau_{x}^{\text{SXPb}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPb}}}\sigma_{x}-\frac{1}{\Omega_{\text{SXPb}}}2g_{r}\left|\alpha_{\text{SXPb}}\right|\sigma_{z},\\ \tau_{y}^{\text{SXPb}}&{}=\sigma_{y},\\ \tau_{z}^{\text{SXPb}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPb}}}\sigma_{z}+\frac{1}{\Omega_{\text{SXPb}}}2g_{r}\left|\alpha_{\text{SXPb}}\right|\sigma_{x}\end{array}$ (42)
for SXPb phase, respectively. The form of unitary operator $U_{k}$ (k=SX, SP, SXPa, SXPb) for k phase reads
$U_{\text{SX}}$ $=\exp[\frac{-i}{\sqrt{2}{\Omega_{\text{SX}}}}\sqrt{\tilde{\omega}_{0}\tilde{\omega}_{c}}(x\tau^{{}_{\text{SX}}}_{y}/\zeta+\zeta^{\prime}p\tau_{x}^{{}_{\text{SX}}})],$ (43)
$U_{\text{SP}}$ $=\exp[\frac{i}{\sqrt{2}{\Omega_{\text{SP}}}}\sqrt{\tilde{\omega}_{0}\tilde{\omega}_{c}}(-\zeta x\tau_{x}^{{}_{\text{SP}}}+p\tau_{y}^{{}_{\text{SP}}}/\zeta^{\prime})],$
$U_{\text{SXPa}}$ $=\exp[\frac{-ig_{c}}{{\sqrt{2}\Omega_{\text{SXPa}}}}(\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPa}}}\tau_{y}^{{}_{\text{SXPa}}}x-\tau_{x}^{{}_{\text{SXPa}}}p)],$
$U_{\text{SXPb}}$ $=\exp[\frac{-ig_{r}}{{\sqrt{2}\Omega_{\text{SXPb}}}}(\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPb}}}\tau_{y}^{{}_{\text{SXPb}}}x+\tau_{x}^{{}_{\text{SXPb}}}p)].$
Appendix B Squeezing parameters
The form of the squeezing parameter $r_{p}$ (p=N, SX, SP, SXPa, SXPb) is given by
$\begin{array}[]{r@{}l}r_{\text{N}}&{}=\frac{1}{4}\ln[(1-\zeta^{2})/(1-\zeta^{\prime^{2}})],\\ r_{\text{SX}}&{}=\frac{1}{4}\ln[(1-\frac{1}{\zeta^{4}})/(1-\frac{\zeta^{\prime 2}}{\zeta^{2}})],\\ r_{\text{SP}}&{}=\frac{1}{4}\ln[(1-\frac{\zeta^{2}}{\zeta^{\prime 2}})/(1-\frac{1}{\zeta^{\prime 4}})],\\ r_{\text{SXPa}}&{}=\infty,\\ r_{\text{SXPb}}&{}=\infty.\end{array}$ (44)

### v2

@@
-Controlling superradiant phase transition in quantum Rabi model
+Controlling the superradiant phase transition in the quantum Rabi model
 Abstract
 In the ultrastrong-coupling regime, the quantum Rabi model can exhibit quantum phase transition (QPT) when the ratio of the qubit transition frequency to the frequency of the cavity field approaches infinity.
-However, it is challenging to control the QPT in few-body systems because of the limited coupling strength and the $A^{2}$ terms.
+However, it is challenging to control the QPT in few-body systems because of the limited coupling strength and the $A^{2}$ term.
 Here, we propose a practical scheme to manipulate the QPT of quantum Rabi model in the strong-coupling regime.
-By applying a periodic frequency modulation to the two-level system in a standard quantum Rabi model in the strong-coupling regime, an anisotropic quantum Rabi model with ultrastrong and tunable coupling strengths for rotating and counter-rotating terms is obtained.
-The ground-state and excitation energy of this model in terms of the modulation parameters are studied.
-We find that the QPT of quantum Rabi model can be observed in the strong-coupling regime and externally controlled by the modulation.
+By applying a periodic frequency modulation to the two-level system in a quantum Rabi model in the strong-coupling regime, an anisotropic quantum Rabi model with ultrastrong and tunable coupling strengths for rotating and counter-rotating terms is obtained.
+The ground state and excitation energy of this model, in terms of the modulation parameters, are studied.
+We find that the QPT of the quantum Rabi model can be observed in the strong-coupling regime and externally controlled by the modulation.
 I Introduction
-Quantum phase transition (QPT) QPT1999 has long been an important and fundamental phenomenon in many-body systems, and has attracted remarkable amount of study interest due to its special features.
+Quantum phase transition (QPT) QPT1999 has long been an important and fundamental phenomenon in many-body systems, and has attracted a remarkable amount of study interest due to its special features.
 The QPT occurs when the non-thermal parameter, such as coupling strength, exceeds a critical value at zero temperature QPT1999 .
 For example, the superradiant phase transition occurs in the Dicke model Dicke when the light-matter coupling strength exceeds a critical value in the thermodynamic limit, i.e., $N\rightarrow\infty$ 1973dickle ; 1973spt ; Sunchangpu ; liyong ; CEPRL ; NLPRL ; ABPRL ; PRL120183603 .
-The control of the QPT in Dicke model has been discussed ABPRL ; 2023Modulation .
-Recently, it is found that the QPT can also occur in few-body systems, such as the quantum Rabi model (QRM) 1936Rabi ; 1937Rabi when the ratio of the qubit transition frequency $\omega_{0}$ to the frequency $\omega_{c}$ of the cavity field tends to infinity, i.e., $\eta=\omega_{0}/\omega_{c}\rightarrow\infty$ Hwang2015Rabi .
-Hereafter, many efforts have been devoted to study the QPT in QRM both theoretically and experimentally trapped ions ; nuclear ; ZhengShiBiao , such as the universal scaling and critical exponents MaoXinLiu2017 , QPT in the quantum Rabi triangle Rabi triangle ; Rabi lattice , dissipative phase transition in the open QRM dispassive QRM ; dar ; sar , QPT in the anisotropic QRM yingzujian ; PRAaqrm and quench dynamics in the anisotropic QRM L.
+The control of the QPT in Dicke model was discussed ABPRL ; 2023Modulation .
+Recently, it was found that the QPT can also occur in few-body systems, such as the quantum Rabi model (QRM) 1936Rabi ; 1937Rabi when the ratio of the qubit transition frequency $\omega_{0}$ to the frequency $\omega_{c}$ of the cavity field tends to infinity, i.e., $\eta=\omega_{0}/\omega_{c}\rightarrow\infty$ Hwang2015Rabi .
+Hereafter, many efforts were devoted to studying the QPT in QRM both theoretically MaoXinLiu2017 ; Rabi triangle ; Rabi lattice ; dispassive QRM ; dar ; sar ; yingzujian ; PRAaqrm ; L.
+T.
+ShenPRA ; full quantum limit and experimentally trapped ions ; nuclear ; ZhengShiBiao , such as the universal scaling and critical exponents MaoXinLiu2017 , QPT in the quantum Rabi triangle Rabi triangle ; Rabi lattice , dissipative phase transition in the open QRM dispassive QRM ; dar ; sar , QPT in the anisotropic QRM yingzujian ; PRAaqrm and quench dynamics in the anisotropic QRM L.
 T.
@@
 Even the QPT in QRM in the fully quantum limit has also been predicted full quantum limit .
-Though the QPT in QRM has been experimentally observed in a single trapped ion trapped ions and a nuclear magnetic resonance quantum simulator nuclear , it is still challenging to manipulate the ground-state QPT due to the limited coupling strength with the classical field (dispersion) limit $\eta\rightarrow\infty$ and the $A^{2}$ terms in cavity quantum electrodynamics (QED) A2 term ; Hwang2015Rabi ; no-go theorm .
-The $A^{2}$ terms can prevent the occurrence of the QPT in QRM Hwang2015Rabi , which can not be neglected in the ultrastrong-coupling regime in cavity QED system.
+Though the QPT in QRM was experimentally observed in a single trapped ion trapped ions and a nuclear magnetic resonance quantum simulator nuclear , it is still challenging to manipulate the ground-state QPT due to the limited coupling strength with the classical field (dispersion) limit $\eta\rightarrow\infty$ and the $A^{2}$ term in cavity quantum electrodynamics (QED) A2 term ; Hwang2015Rabi ; no-go theorm .
+The $A^{2}$ term can prevent the occurrence of the QPT in QRM Hwang2015Rabi , which can not be neglected in the ultrastrong-coupling regime in cavity-QED system.
 In this work, we propose a practical scheme to manipulate the QPT in QRM.
-Specifically, a periodic modulation was applied to the qubit transition frequency in a standard Rabi model 2017manipulation ; C.
+Concretely, a periodic modulation is applied to the qubit transition frequency in a standard QRM 2017manipulation ; C.
 Liu in the strong-coupling regime.
-Then an anisotropic QRM is generated with rotating and counter-rotating coupling strengths allowed to be tuned.
-Through the modulation, both the effective rotating and counter-rotating (CR) coupling strengths can be varied in a wide range and reach ultrastrong-coupling regime.
-The ground state is allowed to be in superradiant phase.
-To investigate the QPT of the anisotropic QRM, we calculated excitation energy and the ground-state phases in both the normal and superradiant phase in the whole coupling parameter space.
+Then we obtain an anisotropic QRM with tunable rotating and counter-rotating coupling strengths.
+Through the modulation, both the effective rotating and counter-rotating (CR) coupling strengths can be varied in a wide range and reach the ultrastrong-coupling regime.
+The ground state is allowed to be in the superradiant phase.
+To investigate the QPT of the anisotropic QRM, we calculate excitation energy and the ground-state phase in both the normal and superradiant phase in the entire coupling parameter space.
 We find that both the normal and superradiant phase can be reached in the strong-coupling regime via the modulation.
 We also find that Goldstone modes Hwang2016JC ; Goldstone exist in the superradiant phase.
-The dependence of excitation energy and ground-state phases on the modulation parameters are analyzed.
+The dependence of excitation energy and ground-state phase on the modulation parameters is analyzed.
 The feature of the QPT is also investigated by calculating the derivatives of the ground-state energy.
@@
 These results may inspire other studies on QPT engineering in other few-body systems.
-This paper is organized as follows.
+The rest of this paper is organized as follows.
 In Sec.
-II, we introduce a periodic modulation to the qubit transition frequency in a standard QRM and obtain an effective anisotropic QRM whose rotationg and counter-rotating terms are tunable.
+II, we introduce a qubit frequency modulation in a standard QRM and obtain an effective anisotropic QRM whose rotating and counter-rotating terms are tunable.
 In Sec.
-III, we analyze the dependence of the effective QPT boundary and coupling strengths on modulation parameters.
+III, we analyze the dependence of the effective frequency ratio and the effective coupling strengths on modulation parameters.
 We obtain the excitation energies for five phases and the phase boundary in Sec.
@@
 In Sec.
-V, we figure out the phase diagram and verify the phase diagram by analyzing the order parameters.
+V, we figure out the phase diagram and verify it by analyzing the order parameters.
 The feature of QPT is also discussed.
-We also discuss how to manipulate the ground-states phases by varying the modulation parameters in Sec.
+We also discuss how to manipulate the ground-state phase by varying the modulation parameters in Sec.
 VI.
 In Sec.
-VII, we discuss the $A^{2}$ terms undsr modulation.
-Finally, we present a brief conclusion and discussions in Sec.
+VII, we discuss the $A^{2}$ term under modulation.
+We present some discussions concerning the experimental implementation in Sec.
 VIII.
+A brief conclusion is given in Sec.
+IX.
+Finally, three Appendixes are presented to show the validity evaluation of the rotating-wave approximation, the operators for superradiant phase, and the squeezing parameters.
+Figure 1: (Color online) Schematic of a cavity field coupled to a two-level system (qubit) with a periodic modulation on the qubit transition frequency.
 II Model and Hamiltonian
-The system we consider is composed by a two-level system (qubit) coupled to a cavity field, as shown in Fig.
+The system we consider is composed of a two-level system (qubit) coupled to a cavity field, as shown in Fig.
 1, which is described by the quantum Rabi model 1936Rabi ; 1937Rabi .
@@
 $H_{\mathrm{\text{JC}}}=\omega_{c}a^{\dagger}a+\frac{\omega_{0}}{2}\sigma_{z}+g(\sigma_{+}a+a^{\dagger}\sigma_{-})$ (2)
-is the usual Jaynes-Cummings (JC) Hamiltonian, and
+is the Jaynes-Cummings (JC) Hamiltonian, and
 $H_{\text{CR}}=g(\sigma_{-}a+a^{\dagger}\sigma_{+})$ (3)
@@
 The transition frequency of the qubit is $\omega_{0}$ and the coupling strength between the qubit and the cavity field is $g$.
+Note that in this work we use the terms of the cavity-QED system to expound the physical idea.
+Actually, the physical results also work for other physical platforms, in which the physical model consisting of a two-level system and a single-mode bosonic field can be implemented.
 To engineer the system, we introduce a sinusoidal modulation to the transition frequency of the qubit. The modulation Hamiltonian is given by 2017manipulation
 $H_{\text{M}}(t)=\frac{1}{2}\xi\nu\cos(\nu t)\sigma_{z},$ (4)
-where $\xi$ and $\nu$ are the dimensionless modulation amplitude and frequency, respectively. After including the two-photon $A^{2}$ terms
+where $\xi$ and $\nu$ are the dimensionless modulation amplitude and frequency, respectively. After including the two-photon $A^{2}$ term
 $H_{A^{2}}=g_{A^{2}}(a+a^{\dagger})^{2}$ (5)
@@
 with $\omega_{c}^{\prime}=\omega_{c}+2g_{A^{2}}$.
-Figure 1: (Color online) Schematic of a cavity field coupled to a two-level system (qubit) with a periodic modulation on the qubit transition frequency.
 In the first rotating frame defined by
-[…]
-where $\mathcal{T}$ denotes the time-ordering operator, the Hamiltonian becomes
+$V_{1}(t)$ $=\exp\left\{-i\left[\left(\omega_{c}^{\prime}a^{\dagger}a+\frac{1}{2}\omega_{0}\sigma_{z}\right)t+\frac{1}{2}\xi\sin(\nu t)\sigma_{z}\right]\right\},$ (8)
+the Hamiltonian becomes
 $\tilde{H}(t)=$ $V_{1}^{\dagger}(t)H(t)V_{1}(t)-iV_{1}^{\dagger}(t)\dot{V}_{1}(t)$
@@
 […]
-where we have omitted the constant term in $H_{A^{2}}$.
+where we omit the constant term in $H_{A^{2}}$.
 Here, $\delta_{n}=\omega_{0}-\omega_{c}^{\prime}+n\nu$ is the oscillating frequency for the $n$th sideband in the rotating terms, and $\Delta_{m}=\omega_{0}+\omega_{c}^{\prime}+m\nu$ is the oscillating frequency for the $m$th sideband in the CR terms.
-In the derivation of $\tilde{H}(t)$, we have used the Jacobi-Anger expansion
-$e^{i\xi\sin(\nu t)}=\sum_{n=-\infty}^{\infty}J_{n}(\xi)e^{in\nu t}.$ (10)
-Here, $J_{n}(\xi)$ is the $n$th Bessel function of the first kind.
+In the derivation of $\tilde{H}(t)$, we use the Jacobi-Anger expansion
+$e^{i\xi\sin(\nu t)}=\sum_{n=-\infty}^{\infty}J_{n}(\xi)e^{in\nu t},$ (10)
+where $J_{n}(\xi)$ is the $n$th Bessel function of the first kind.
 In Eq.
@@
 The oscillating frequencies of all rotating and CR sidebands can be written as $\delta_{n_{0}+s}=\delta_{n_{0}}+s\nu$ and $\Delta_{m_{0}+s}=\Delta_{m_{0}}+s\nu$ with $s$ an integer, respectively.
-Thus the other sidebands are separated from $n_{0}$ ($m_{0}$) sideband by $s\nu$.
+Thus the other sidebands are separated from the $n_{0}$ ($m_{0}$) sideband by $s\nu$.
 Under the conditions
 $\nu\gg g>g\left|J_{m}(\xi)\right|,\ g\left|J_{n}(\xi)\right|\ \text{and}\ \nu\gg\left|\Delta_{m_{0}}\right|,\ \left|\delta_{n_{0}}\right|,$ (11)
-all the sidebands other than the $n_{0}$($m_{0}$) sideband can be discarded by using RWA.
-With $g_{A^{2}}\ll g\ll\omega_{0}$ and $g_{A^{2}}\ll 2\omega_{c}^{\prime}$, the $a^{2}$ and $a^{\dagger 2}$ terms with oscillating frequencies ±$\pm 2\omega_{c}^{\prime}$ from the two-photon $A^{2}$ Hamiltonian can also be neglected using RWA.
-To be focus, the validity of this approximation is discussed in detail in Sec.
+all the sidebands other than the $n_{0}$($m_{0}$)th sideband can be discarded using the rotating-wave approximation (RWA).
+With $g_{A^{2}}\ll g\ll\omega_{0}$ and $g_{A^{2}}\ll 2\omega_{c}^{\prime}$, the $a^{2}$ and $a^{\dagger 2}$ terms with oscillating frequencies ±$\pm 2\omega_{c}^{\prime}$ from the two-photon $A^{2}$ Hamiltonian can also be neglected using the RWA.
+To be precise, the validity of this approximation is discussed in detail in Sec.
 VII.
@@
 […]
-where we have introduced the effective coupling strengths
-$g_{r}(\xi)=gJ_{n_{0}}(\xi),\ g_{cr}(\xi)=gJ_{m_{0}}(\xi).$ (13)
-The Hamiltonian describes an effective standard Rabi model with $g_{r}=g_{cr}$ or an effective anisotropic QRM with $g_{r}\neq g_{cr}$ QTxie in interaction picture.
+where we introduce the effective coupling strengths
+$g_{r}(\xi)=gJ_{n_{0}}(\xi),\hskip 14.22636ptg_{cr}(\xi)=gJ_{m_{0}}(\xi).$ (13)
+To evaluate the validity of the RWA made in Hamiltonian (9), we examine the dynamics of the fidelity between the evolving states governed by Hamiltonians (9) and (12).
+The results are shown in Appendix A.
+The Hamiltonian describes an effective standard QRM with $g_{r}=g_{cr}$ or an effective anisotropic QRM with $g_{r}\neq g_{cr}$ QTxie in interaction picture.
 Note that, we can tune the effective rotating and CR coupling strengths $g_{r}(\xi)$ and $g_{cr}(\xi)$ via the dimensionless modulation amplitude $\xi$.
-To make Hamiltonian (12) time independent, we introduce the effective frequency for the qubit
-$\tilde{\omega}_{0}$ $=\frac{\Delta_{m_{0}}+\delta_{n_{0}}}{2},$
-and the cavity field
-$\tilde{\omega}_{c}=\frac{\Delta_{m_{0}}-\delta_{n_{0}}}{2}.$
+To make Hamiltonian (12) time independent, we introduce the effective frequencies
+[…]
+for the qubit and the cavity field, respectively.
+Here, $\tilde{\omega}_{0}(\nu)$ and $\tilde{\omega}_{c}(\nu)$ depend on the modulation frequency $\nu$ via $\Delta_{m_{0}}$ and $\delta_{n_{0}}$.
+Hereafter, we denote $\tilde{\omega}_{0,c}(\nu)$ as $\tilde{\omega}_{0,c}$ for precision.
 After transferring to the second rotating frame by the unitary transformation
@@
 […]
-Here, we have introduced the anisotropic parameter $\epsilon(\xi)=g_{cr}(\xi)/g_{r}(\xi)$.
-The effective coupling strength $g_{r}(\xi)$, the effective frequencies $\tilde{\omega}_{0}$ and $\tilde{\omega}_{c}$, and anisotropic parameter $\epsilon(\xi)$ can be regulated by the modulation parameters $\nu$ and $\xi$.
+Here, we introduce the anisotropic parameter $\epsilon(\xi)=g_{cr}(\xi)/g_{r}(\xi)$.
+The effective coupling strength $g_{r}(\xi)$, the effective frequencies $\tilde{\omega}_{0}(\nu)$ and $\tilde{\omega}_{c}(\nu)$, and the anisotropic parameter $\epsilon(\xi)$ can be tuned by the modulation parameters $\nu$ and $\xi$.
 The effective Hamiltonian (16) describe an anisotropic QRM with $\epsilon(\xi)\neq 1$ and a standard QRM with $\epsilon(\xi)=1$.
-Hereafter, we write $\epsilon(\xi)$ as $\epsilon$ for short.
+Hereafter, we denote $\epsilon(\xi)$ as $\epsilon$ for precision.
+Figure 2: (a) The effective ratio $\tilde{\eta}$ versus the modulation frequency $\nu/\omega_{0}$ for $\eta=10^{2}\text{ and }10^{3}.$ (b), (c) The couplings $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ versus the modulation amplitude $\xi$.
+Here, $\nu/\omega_{0}=0.49$ and $\eta=1$ in panel (b) and $\nu/\omega_{0}=0.68$ and $\eta=10^{2}$ in panel (c), the other parameter is $g/\omega_{0}=0.06.$
 III Ultrastrong coupling
-The effective standard quantum Rabi model ($\epsilon=1$) in the limit $\tilde{\eta}=\tilde{\omega}_{0}/\tilde{\omega}_{c}\rightarrow\infty$ can exhibit quantum phase transition from the normal phase to the superradiant phase.
+The effective QRM ($\epsilon=1$) in the limit $\tilde{\eta}=\tilde{\omega}_{0}/\tilde{\omega}_{c}\rightarrow\infty$ can exhibit quantum phase transition from the normal phase to the superradiant phase.
 This phase transition occurs at $g_{C}=\sqrt{\tilde{\omega}_{0}\tilde{\omega}_{c}}/2$ Ashhab2013 above which $Z_{2}$ parity symmetry is broken and the bosonic mode is macroscopically occupied Hwang2015Rabi .
@@
 As shown in Fig.
-2(a), the curve of the effective ratio $\tilde{\eta}$ is composed by series of V-type valleys until $\nu$ increases up to $\nu/\omega_{0}=2(\omega_{0}+\omega_{c}^{\prime})$.
+2(a), the curve of the effective ratio $\tilde{\eta}$ is composed of a series of V-type valleys until $\nu$ increases up to $\nu/\omega_{0}=2(\omega_{0}+\omega_{c}^{\prime})$.
 The V-shaped valleys share the same $m_{0}$ and the valley for $m_{0}$ starts at $\nu=2(\omega_{0}+\omega_{c}^{\prime})/(2\left|m_{0}\right|+1)$.
@@
 Thus, in the regime $\nu<2(\omega_{0}+\omega_{c}^{\prime})$, we can manipulate $\tilde{\eta}$ via choosing appropriate $\nu$ while in the regime $\nu>2(\omega_{0}+\omega_{c}^{\prime})$, this $\tilde{\eta}$ cannot be varied via changing $\nu$.
+As shown in Fig.
+2(a), to satisfy the condition (11) and the limit $\tilde{\eta}\rightarrow\infty$, the value of $\nu$ can be chosen in a broad range.
 To see clearly the dependence of the effective coupling strengths $g_{r}(\xi)$ and $g_{cr}(\xi)$ on the modulation amplitude $\xi$, we plot $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ versus the modulation amplitude $\xi$ for $\eta=1$ and $10^{2}$ at $\nu/\omega_{0}=0.49$ and $0.68$ in Figs.
@@
 The effective coupling strength $g_{r}/g_{C}\ (g_{cr}/g_{C})$ depends on the modulation amplitude $\xi$ in the form of the Bessel function.
-When $\nu/\omega_{0}=0.49$ for $\eta=1$ $(m_{0}=-4,\ n_{0}=-1)$, $g_{r}/g_{C}$ oscillates between -2.41656 and 6 with reducing amplitude, and $g_{cr}/g_{C}$ oscillates between -1.64629 and 2.39791, as shown in Fig.
+When $\nu/\omega_{0}=0.49$ for $\eta=1$ $(m_{0}=-4,\ n_{0}=-1)$, $g_{r}/g_{C}$ oscillates between $-2.41656$ and 6 with reducing amplitude, and $g_{cr}/g_{C}$ oscillates between $-1.646$ and 2.398, as shown in Fig.
 2(b).
-When $\nu/\omega_{0}=0.68$ for $\eta=10^{2}$ $(m_{0}=n_{0}=-1)$, $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ show the same oscillating feature between -1.23432 and 0.734245, as shown in Fig.
+When $\nu/\omega_{0}=0.68$ for $\eta=10^{2}$ $(m_{0}=n_{0}=-1)$, $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ show the same oscillating feature between $-1.234$ and 0.734, as shown in Fig.
 2(c).
-Note that, $g_{r}=g_{cr}$ in Fig.
-2(c) under the chosen parameters.
 The results show that one can tune the effective coupling strengths $g_{r}$ and $g_{cr}$ in a broad range.
-Figure 2: (Color online) (a) The effective ratio $\tilde{\eta}$ versus the modulation frequency $\nu/\omega_{0}$ for $\eta=10^{2}\text{ and }10^{3}.$ (b-c) The couplings $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ versus the modulation amplitude $\xi$.
-Here, $\nu/\omega_{0}=0.49$ and $\eta=1$ in panel (b) and $\nu/\omega_{0}=0.68$ and $\eta=10^{2}$ in panel (c), other parameter is $g/\omega_{0}=0.06.$
+Figure 3: The phase transition boundary $\tilde{g}_{C}/\omega_{0}$ versus $\xi$ for the anisotropic QRM at $\eta=10^{2}$ and $\nu/\omega_{0}=0.402$, $0.68$, and $0.8$.
 IV Excitation spectra
-To figure out the ground-state property of the system in the whole coupling-parameter space, we calculate the excitation energy in the limit $\tilde{\eta}\rightarrow\infty$ in this section.
+To figure out the ground-state property of the system in the entire coupling-parameter space, we calculate the excitation energy in the limit $\tilde{\eta}\rightarrow\infty$ in this section.
 IV.1 Normal phase
@@
 $\tilde{H}_{\text{I}}$ $=\frac{g_{r}}{\sqrt{2}}[(1+\epsilon)\sigma_{x}x-(1-\epsilon)\sigma_{y}p].$ (18b)
-Then, in the limit $\tilde{\eta}\rightarrow\infty$, we make an unitary transformation $U_{\text{N}}^{\dagger}{H}_{\text{AR}}U_{\text{N}}$ on Hamiltonian (16) with
+Then, in the limit $\tilde{\eta}\rightarrow\infty$, we make a unitary transformation $U_{\text{N}}^{\dagger}{H}_{\text{AR}}U_{\text{N}}$ on Hamiltonian (17) with
 $U_{\text{N}}=\exp\{-ig_{r}[(1-\epsilon)\sigma_{x}p+(1+\epsilon)\sigma_{y}x]/{\sqrt{2}\tilde{\omega}_{0}}\}.$ (19)
 The transformed Hamiltonian becomes free of coupling between qubit subspaces $\mathcal{H}_{g}$ and $\mathcal{H}_{e}$, which are spanned by $\left|g\right\rangle$ and $\left|e\right\rangle$, respectively.
-After a projection into $\mathcal{H}_{g}$, i.e., $\tilde{H}_{\text{N}}=\langle g|U_{\text{N}}^{\dagger}H_{\text{AR}}U_{\text{N}}|g\rangle$, we obtain an effective low-energy Hamiltonian as
+After a projection into $\mathcal{H}_{g}$, i.e., $\tilde{H}_{\text{N}}=\langle g|U_{\text{N}}^{\dagger}H_{\text{AR}}U_{\text{N}}|g\rangle$, we obtain an effective low-energy Hamiltonian
 $\tilde{H}_{\text{N}}=\frac{{\tilde{\omega}_{c}}}{2}[(1-\zeta^{2})x^{2}+(1-\zeta^{\prime^{2}})p^{2}],$ (20)
-where a constant term is neglected. Here,
-Figure 3: (Color online) The phase transition boundary $\tilde{g}_{C}/\omega_{0}$ for the anisotropic QRM versus $\xi$ for $\nu/\omega_{0}=0.402$ and $0.68$ with $\eta=10^{2}$.
-$\zeta=\frac{\lambda(1+\epsilon)}{2}\ \text{and}\ \zeta^{\prime}=\frac{\lambda(1-\epsilon)}{2}$ (20)
+upto a constant term, where
+$\zeta=\frac{\lambda(1+\epsilon)}{2},\hskip 14.22636pt\zeta^{\prime}=\frac{\lambda(1-\epsilon)}{2},$ (21)
 with the dimensionless rotating-coupling strength $\lambda\equiv{g_{r}/g_{C}}$.
@@
 (20).
-The excitation energy $\omega_{\text{N}}$ can be directly obtained as
+From Eq.
+(20), The excitation energy $\omega_{\text{N}}$ can be directly obtained as
 $\omega_{\text{N}}=\tilde{\omega}_{c}\sqrt{(1-\zeta^{2})(1-\zeta^{\prime 2})}.$ (22)
 $\tilde{H}_{\text{N}}$ implies divergent values of $\langle x\rangle$ ($\langle p\rangle$) due to the disappearance of $x^{2}$ ($p^{2}$) term when $\epsilon>0$ ($\epsilon<0$) in the limit $\tilde{\eta}\rightarrow\infty$.
-By this fact, the normal phase is restricted in the region
+Due to this fact, the normal phase is restricted in the region
 $g_{r}<\left|\tilde{g}_{C}\right|$ (23)
@@
 Note that, $g_{C}$ is the critical coupling strength for the standard QRM.
-When $g_{r}<\left|\tilde{g}_{C}\right|$, the ground state is in the normal phase with the cavity field in vacuum and qubit in $\left|g\right\rangle$.
-While outside the normal phase, namely, $g_{r}>\left|\tilde{g}_{C}\right|$, the system is in the superradiant phase with both the qubit and the cavity field excited.
+When $g_{r}<\left|\tilde{g}_{C}\right|$, the ground state is in the normal phase with the cavity field in a vacuum and qubit in $\left|g\right\rangle$.
+While outside the normal phase, namely, $g_{r}>\left|\tilde{g}_{C}\right|$, the system is in the superradiant phase with both the qubit and the cavity field being excited.
 In Fig.
-3, we plot $\tilde{g}_{C}/\omega_{0}$ versus $\xi$ for $\nu/\omega_{0}=0.402\ \text{and}\ 0.68$.
-For $\nu/\omega_{0}=0.402$, $\tilde{g}_{C}/\omega_{0}$ oscillates between 0 and 0.0309031.
-The zero points of $\tilde{g}_{C}/\omega_{0}$ locate at those zeros of $J_{n_{0}}(\xi)$.
-The first zero point of $\tilde{g}_{C}/\omega_{0}$ locates at $\xi=5.13562$.
-For a larger $\nu$, such as $\nu/\omega_{0}=0.68$, $\tilde{g}_{C}/\omega_{0}$ becomes independent of $\xi$ due to $m_{0}=n_{0}$.
-Thus, we can tune the QPT boundary via $\nu$ and $\xi$.
+3, we plot $\tilde{g}_{C}/\omega_{0}$ versus $\xi$ for $\nu/\omega_{0}=0.402$, $0.68$, and $0.8$.
+For $\nu/\omega_{0}=0.402$, $\tilde{g}_{C}/\omega_{0}$ oscillates between 0 and 0.031.
+The zero points of $\tilde{g}_{C}/\omega_{0}$ are located at $\xi=5.136$, $8.417$, and $11.62$, corresponding to the zeros of $J_{n_{0}}(\xi)$ ($n_{0}=-2$).
+For $(1+1/\eta)/(\left|n\right|+3/2)\leqslant\nu/\omega_{0}<(1-1/\eta)/(\left|n\right|+1/2)$ with $n\in\left\{-N\right\}$ and $\nu/\omega_{0}\geq 2(1+1/\eta)$, $\tilde{g}_{C}/\omega_{0}$ becomes independent of $\xi$ due to $m_{0}=n_{0}$.
+For example, when $\eta=10^{2}$ and $n=0$, the value of $\nu/\omega_{0}$ is between 0.673 and 1.98, such as $\nu/\omega_{0}=0.68$ and $0.8$ shown in Fig.
+3.
+When $m_{0}=n_{0}$, the QPT boundary $\tilde{g}_{C}$ of the anisotropic QRM returns back to the critical point ${g}_{C}$ of the QRM.
+Thus, the QPT boundary can be tuned via $\nu$ and $\xi$.
 IV.2 Superradiant phase
 Now, we study the excitation energy in the superradiant phase, namely, $g_{r}>\left|\tilde{g}_{C}\right|$.
-In the superradiant phase, there are four different superradiant phases.
+In the superradiant phase, there are four different phases.
 When $\epsilon>0$, the superradiant phase can be defined as superradiant $X$ (SX) phase since the $\langle x\rangle$ becomes divergent.
@@
 For $g_{cr}>2\left|g_{C}\right|$ with $g_{r}=0$, the ground state of system is in superradiant XP $a$ (SXPa) phase while for $g_{r}>2\left|g_{C}\right|$ with $g_{cr}=0$, the system is in superradiant XP $b$ (SXPb) phase.
-When $g_{cr}=0$, only rotating terms exist in the Hamiltonian, which reduces to a Jaynes-Cummings model.
-In the following, we investigate the excitation energy in the limit $\tilde{\eta}\rightarrow\infty$ for SX, SP, SXPa and SXPb phase, respectively.
-To effectively describe the superradiant phase with macroscopic excitation, we displace the cavity field $a$ by displacing operator $D_{k}(\alpha)$ (k=SX, SP, SXPa, SXPb) for k phase.
+When $g_{cr}=0$, only rotating terms exist and the Hamiltonian reduces to a Jaynes-Cummings model.
+In the following, we investigate the excitation energy in the limit $\tilde{\eta}\rightarrow\infty$ for SX, SP, SXPa, and SXPb phase, respectively.
+To effectively describe the superradiant phase with macroscopic excitation, we displace the cavity field $a$ by the displacing operator $D_{k}(\alpha)$ ($k$=SX, SP, SXPa, SXPb) for the $k$ phase.
 After the displacement, the Hamiltonian $H_{\text{AR}}$ becomes
 $\tilde{H}_{AR}=D_{k}^{\dagger}(\alpha)H_{\text{AR}}D_{k}(\alpha).$ (25)
-Here, the detailed form of $D_{k}(\alpha)$ is given in the Appendix A.
+Here, the detailed form of $D_{k}(\alpha)$ is given in the Appendix B.
 To diagonalize the qubit part in Hamiltonian $\tilde{H}_{AR}$, we introduce a new set of Pauli operators $\tau_{x,y,z}^{k}$ ($k$=SX, SP, SXPa, SXPb) for each k phase, respectively.
-To be concise, the form of $\tau_{x,y,z}^{k}$ for k phase is also given in the Appendix A.
+To be concise, the form of $\tau_{x,y,z}^{k}$ for the $k$ phase is also given in the Appendix B.
 In terms of the new Pauli operators $\tau_{x,y,z}^{k}$ , the displaced Hamiltonian $\tilde{H}_{\text{AR}}$ becomes
@@
 $+\frac{\tilde{\omega}_{c}}{2}\alpha_{\text{SP}}^{2},$
-$\tilde{H}_{\text{AR}}(\alpha_{\text{SXPa}})=$ $\frac{\Omega_{\text{SXPa}}}{2}\tau_{z}^{{}_{\text{SXPa}}}+\chi_{3}(\frac{4}{\mu}\tau_{x}^{{}_{\text{SXPa}}}x+\mu\tau_{y}^{{}_{\text{SXPa}}}p)$
+$\tilde{H}_{\text{AR}}(\alpha_{\text{SXPa}})=$ $\frac{\Omega_{\text{SXPa}}}{2}\tau_{z}^{{}_{\text{SXPa}}}+\chi_{3}\left(\frac{4}{\mu}\tau_{x}^{{}_{\text{SXPa}}}x+\mu\tau_{y}^{{}_{\text{SXPa}}}p\right)$
 $+H_{c}+\tilde{\omega}_{c}\alpha_{\text{SXPa}}^{2},$
-$\tilde{H}_{\text{AR}}(\alpha_{\text{SXPb}})=$ $\frac{\Omega_{\text{SXPb}}}{2}\tau_{z}^{{}_{\text{SXPb}}}+\chi_{3}(\frac{4}{\lambda}\tau_{x}^{{}_{\text{SXPb}}}x-\lambda\tau_{y}^{{}_{\text{SXPb}}}p)$
+$\tilde{H}_{\text{AR}}(\alpha_{\text{SXPb}})=$ $\frac{\Omega_{\text{SXPb}}}{2}\tau_{z}^{{}_{\text{SXPb}}}+\chi_{3}\left(\frac{4}{\lambda}\tau_{x}^{{}_{\text{SXPb}}}x-\lambda\tau_{y}^{{}_{\text{SXPb}}}p\right)$
 $+H_{c}+\tilde{\omega}_{c}\alpha_{\text{SXPb}}^{2}.$
@@
 $H_{c}=\frac{\tilde{\omega}_{c}}{2}(x^{2}+p^{2})-\frac{\tilde{\omega}_{c}}{2},$ (27)
-and the rescaled qubit frequency $\Omega_{k}$ for k phase is
+and the rescaled qubit frequency $\Omega_{k}$ for the $k$ phase is
 $\begin{array}[]{r@{}l}\Omega_{\text{SX}}&{}=\tilde{\omega}_{0}\zeta^{2},\\ \Omega_{\text{SP}}&{}=\tilde{\omega}_{0}\zeta^{\prime 2},\\ \Omega_{\text{SXPa}}&{}=g_{cr}^{2}/\tilde{\omega}_{c},\\ \Omega_{\text{SXPb}}&{}=g_{r}^{2}/\tilde{\omega}_{c}.\end{array}$ (28)
-The displacement parameter $\alpha$ is determined by the vanishment of the non-diagonal terms during deriving Eq. (25), which is obtained as
-$\alpha_{\text{SX}}$ =±$=\pm\sqrt{\frac{\tilde{\eta}}{2}(\zeta^{2}-\frac{1}{\zeta^{2}})},$ (28)
-$\alpha_{\text{SP}}$ =±$=\pm\sqrt{\frac{\tilde{\eta}}{2}(\zeta^{\prime 2}-\frac{1}{\zeta^{\prime 2}})},$
-$\alpha_{\text{SXPa}}$ =±$=\pm\sqrt{\tilde{\eta}(\frac{\mu^{2}}{16}-\frac{1}{\mu^{2}})},$
-$\alpha_{\text{SXPb}}$ =±$=\pm\sqrt{\tilde{\eta}(\frac{\lambda^{2}}{16}-\frac{1}{\lambda^{2}})},$
-for k (k=SX, SP, SXPa, SXPb) phase.
-To distinguish $\alpha$ in each k phase, here we have added a subscript “k” in $\alpha$, i.e., we label $\alpha$ as $\alpha_{k}$.
+The displacement parameter $\alpha$ is determined by the vanishing of the non-diagonal terms during deriving Eq. (26), which is obtained as
+$\alpha_{\text{SX}}$ =±$=\pm\sqrt{\frac{\tilde{\eta}}{2}\left(\zeta^{2}-\frac{1}{\zeta^{2}}\right)},$ (29)
+$\alpha_{\text{SP}}$ =±$=\pm\sqrt{\frac{\tilde{\eta}}{2}\left(\zeta^{\prime 2}-\frac{1}{\zeta^{\prime 2}}\right)},$
+$\alpha_{\text{SXPa}}$ =±$=\pm\sqrt{\tilde{\eta}\left(\frac{\mu^{2}}{16}-\frac{1}{\mu^{2}}\right)},$
+$\alpha_{\text{SXPb}}$ =±$=\pm\sqrt{\tilde{\eta}\left(\frac{\lambda^{2}}{16}-\frac{1}{\lambda^{2}}\right)},$
+for the $k$ ($k$=SX, SP, SXPa, SXPb) phase.
+To distinguish $\alpha$ in each $k$ phase, here we add a subscript $k$ in $\alpha$, i.e., we label $\alpha$ as $\alpha_{k}$.
 Here, the dimensionless CR coupling strength $\mu$ is defined as $\mu=g_{cr}/g_{C}$.
@@
 […]
-Thus, we can make similar unitary transformation $U_{k}$ as $U_{\text{N}}$ in the normal phase on $\tilde{H}_{\text{AR}}(\alpha_{k})$ to eliminate the non-diagonal terms, namely,
+Thus, we can make a similar unitary transformation $U_{k}$ as $U_{\text{N}}$ in the normal phase on $\tilde{H}_{\text{AR}}(\alpha_{k})$ to eliminate the non-diagonal terms, namely,
 $\tilde{H}_{k}=U_{k}^{\dagger}\tilde{H}_{\text{AR}}(\alpha_{k})U_{k},$ (31)
 ($k$=SX, SP, SXPa, SXPb).
-Here, the detailed form of $U_{k}$ is given in the Appendix A.
-After performing the similar procedure used to derive $\tilde{H}_{\text{N}}$, we obtain the effective Hamiltonian $\tilde{H}_{k}$ in k phase as following:
-$\tilde{H}_{\text{SX}}$ $=\frac{\tilde{\omega}_{c}}{2}(1-\frac{1}{\zeta^{4}})x^{2}+\frac{\tilde{\omega}_{c}}{2}(1-\frac{\zeta^{\prime 2}}{\zeta^{2}})p^{2},$ (31)
-$\tilde{H}_{\text{SP}}$ $=\frac{\tilde{\omega}_{c}}{2}(1-\frac{\zeta^{2}}{\zeta^{\prime 2}})x^{2}+\frac{\tilde{\omega}_{c}}{2}(1-\frac{1}{\zeta^{\prime 4}})p^{2},$
-$\tilde{H}_{\text{SXPa}}$ $=\frac{\tilde{\omega}_{c}}{4}(1-\frac{16}{\mu^{4}})x^{2}-\frac{\tilde{\omega}_{c}}{2}(1+\frac{4}{\mu^{2}})-\tilde{\omega}_{0}(\frac{\mu^{2}}{16}+\frac{1}{\mu^{2}}),$
-$\tilde{H}_{\text{SXPb}}$ $=\frac{\tilde{\omega}_{c}}{4}(1-\frac{16}{\lambda^{4}})x^{2}-\frac{\tilde{\omega}_{c}}{2}(1-\frac{4}{\lambda^{2}})-\tilde{\omega}_{0}(\frac{\lambda^{2}}{16}+\frac{1}{\lambda^{2}}).$
+Here, the detailed form of $U_{k}$ is given in the Appendix B.
+After performing a similar procedure used to derive $\tilde{H}_{\text{N}}$, we obtain the effective Hamiltonian $\tilde{H}_{k}$ in the $k$ phase as follows:
+$\tilde{H}_{\text{SX}}$ $=\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{1}{\zeta^{4}}\right)x^{2}+\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{\zeta^{\prime 2}}{\zeta^{2}}\right)p^{2},$ (32)
+$\tilde{H}_{\text{SP}}$ $=\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{\zeta^{2}}{\zeta^{\prime 2}}\right)x^{2}+\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{1}{\zeta^{\prime 4}}\right)p^{2},$
+$\tilde{H}_{\text{SXPa}}$ $=\frac{\tilde{\omega}_{c}}{4}\left(1-\frac{16}{\mu^{4}}\right)x^{2}-\frac{\tilde{\omega}_{c}}{2}\left(1+\frac{4}{\mu^{2}}\right)-\tilde{\omega}_{0}\left(\frac{\mu^{2}}{16}+\frac{1}{\mu^{2}}\right),$
+$\tilde{H}_{\text{SXPb}}$ $=\frac{\tilde{\omega}_{c}}{4}\left(1-\frac{16}{\lambda^{4}}\right)x^{2}-\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{4}{\lambda^{2}}\right)-\tilde{\omega}_{0}\left(\frac{\lambda^{2}}{16}+\frac{1}{\lambda^{2}}\right).$
 Note that Hamiltonian $\tilde{H}_{k}$ in Eq.
@@
 Now, the excitation energy $\omega_{k}$ for each $k$ phase can be obtained as
-$\begin{array}[]{r@{}l}\omega_{\text{SX}}&{}=\tilde{\omega}_{c}\sqrt{(1-\frac{1}{\zeta^{4}})(1-\frac{\zeta^{\prime 2}}{\zeta^{2}})},\\ \omega_{\text{SP}}&{}=\tilde{\omega}_{c}\sqrt{(1-\frac{1}{\zeta^{\prime 4}})(1-\frac{\zeta^{2}}{\zeta^{\prime 2}})},\\ \omega_{\text{SXPa}}&{}=0,\\ \omega_{\text{SXPb}}&{}=0.\end{array}$ (32)
-We find that the excitation energy closes both in SXPa and SXPb phases.
-This result denotes the existence of the Goldstone modes Hwang2016JC ; Goldstone .
-Furthermore, $\omega_{\text{{SP}}}$ has the same form as $\omega_{\text{{SX}}}$ if making the replacement $\zeta\leftrightarrow$$\zeta^{\prime}$ in $\omega_{\text{{SX}}}$.
-[…]
-V Phase diagram
-In this section, we investigate the phase diagram based on the above analysis and order parameters.
-Then we discuss the feature of QPTs by studying the discontinuity of the derivatives of the ground-state energy $E_{G}$.
-Figure 5: (Color online) $\langle x\rangle$ in panel (a) and $\langle p\rangle$ in panel (b) versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ for $\eta=10^{2}$.
-To study the phase diagram, we introduce the average values $\langle x\rangle$ and $\langle p\rangle$ as the order parameters.
-The average values $\langle x\rangle$ and $\langle p\rangle$, and the corresponding variances $\Delta x^{2}$ and $\Delta p^{2}$ for p phase are obtained from the ground state of Hamiltonian $\tilde{H}_{p}$ (p=N, SX, SP, SXPa, SXPb).
-We show the results in table 1.
-The expressions of $s$, $v$ and $w$ in table 1 are defined as
-[…]
-$v(x_{1},x_{2})$ $=\frac{1}{2}\sqrt{\frac{1-x_{2}^{2}}{1-x_{1}^{2}}},$ (33b)
-$w(x_{1},x_{2})$ $=\frac{\sqrt{x_{1}}}{1+x_{1}}\frac{x_{2}^{2}}{\sqrt{x_{2}^{4}-1}}.$ (33c)
-We find that, in the limit $\tilde{\eta}\rightarrow\infty$, $\langle x\rangle\rightarrow\infty$ and $\langle p\rangle=0$ in SX phase while $\langle x\rangle=0$ and $\langle p\rangle\rightarrow\infty$ in the SP phase.
-That is why they are defined as SX phase and SP phase, respectively.
-We also note that, the variances $\Delta x^{2}$ and $\Delta p^{2}$ are finite for N, SX and SP phases, but $\Delta x^{2}=0$ and $\Delta p^{2}=\infty$ for both SXPa and SXPb phases.
-Table 1: $\langle x\rangle$, $\langle p\rangle$, $\langle x^{2}\rangle$ and $\langle p^{2}\rangle$ for all phases.
+$\begin{array}[]{r@{}l}\omega_{\text{SX}}&{}=\tilde{\omega}_{c}\sqrt{\left(1-\frac{1}{\zeta^{4}}\right)\left(1-\frac{\zeta^{\prime 2}}{\zeta^{2}}\right)},\\ \omega_{\text{SP}}&{}=\tilde{\omega}_{c}\sqrt{\left(1-\frac{1}{\zeta^{\prime 4}}\right)\left(1-\frac{\zeta^{2}}{\zeta^{\prime 2}}\right)},\\ \omega_{\text{SXPa}}&{}=0,\\ \omega_{\text{SXPb}}&{}=0.\end{array}$ (33)
+Table 1: $\langle x\rangle$, $\langle p\rangle$, $\langle x^{2}\rangle$, and $\langle p^{2}\rangle$ for all phases.
 N SX SP SXPa SXPb
@@
 $\Delta p^{2}$ $v(\zeta^{\prime},\zeta)$ $4^{-1}w^{-1}(\epsilon,\zeta)$ $w(-\epsilon,\zeta^{\prime})$infty $\infty$
+We find that the excitation energy closes both in the SXPa and SXPb phases.
+This result denotes the existence of the Goldstone modes Hwang2016JC ; Goldstone .
+Furthermore, $\omega_{\text{{SP}}}$ has the same form as $\omega_{\text{{SX}}}$ if making the replacement $\zeta\leftrightarrow$$\zeta^{\prime}$ in $\omega_{\text{{SX}}}$.
+Figure 5: $\langle x\rangle$ in panel (a) and $\langle p\rangle$ in panel (b) versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ for $\eta=10^{2}$.
+[…]
+Figure 6: (a), (b) The first derivatives $\partial E_{G}/\partial\lambda$ and $\partial E_{G}/\partial\mu$ versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$.
+(c), (d) The second derivatives $\partial^{2}E_{G}/\partial\lambda^{2}$ and $\partial^{2}E_{G}/\partial\mu^{2}$ versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$.
+Other parameters are $\omega_{0}=1,\eta=10^{2}$.
+V Phase diagram
+In this section, we investigate the phase diagram based on the above analysis and order parameters.
+Then we discuss the feature of QPTs by studying the discontinuity of the derivatives of the ground-state energy $E_{G}$.
+To study the phase diagram, we introduce the average values $\langle x\rangle$ and $\langle p\rangle$ as the order parameters.
+The average values $\langle x\rangle$ and $\langle p\rangle$, and the corresponding variances $\Delta x^{2}$ and $\Delta p^{2}$ for $p$ phase are obtained from the ground state of the Hamiltonian $\tilde{H}_{p}$ ($p$=N, SX, SP, SXPa, SXPb).
+We show the results in Table 1.
+The expressions of $s$, $v$, and $w$ in Table 1 are defined as
+Figure 7: (Color online) The excitation energy $\omega$ versus the dimensionless modulation amplitude $\xi$. Parameters are $\eta=10^{2},\nu/\omega_{0}=0.68$, and $g/\omega_{0}=0.06$.
+[…]
+$v(x_{1},x_{2})$ $=\frac{1}{2}\sqrt{\frac{1-x_{2}^{2}}{1-x_{1}^{2}}},$ (34b)
+$w(x_{1},x_{2})$ $=\frac{\sqrt{x_{1}}}{1+x_{1}}\frac{x_{2}^{2}}{\sqrt{x_{2}^{4}-1}}.$ (34c)
+Figure 8: (Color online) (a) The parameterized curve of the couplings $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ with the modulation amplitude $\xi$ increasing from 0 to 3 as indicated by the arrows.
+(b) $\langle x\rangle$ versus the modulation amplitude $\xi$.
+The parameters are the same as Fig.
+7.
+It is found that, in the limit $\tilde{\eta}\rightarrow\infty$, we have $\langle x\rangle\rightarrow\infty$ and $\langle p\rangle=0$ in the SX phase while $\langle x\rangle=0$ and $\langle p\rangle\rightarrow\infty$ in the the SP phase.
+That is why they are defined as the SX phase and SP phase, respectively.
+We also note that the variances $\Delta x^{2}$ and $\Delta p^{2}$ are finite for N, SX, and SP phases, but $\Delta x^{2}=0$ and $\Delta p^{2}=\infty$ for both the SXPa and SXPb phases.
 To see clearly the dependence of $\langle x\rangle$ and $\langle p\rangle$ on the coupling strengths $g_{r}$ and $g_{cr}$, we plot $\langle x\rangle$ and $\langle p\rangle$ versus the coupling strengths $g_{r}$ and $g_{cr}$ in Fig.
 5.
-In the region $g_{r}<\left|\tilde{g}_{C}\right|$, both $\langle x\rangle$ and $\langle p\rangle$ are 0, which correspond to the normal phase.
-Outside the normal phase region, i.e., $g_{r}>\left|\tilde{g}_{C}\right|$, when $\epsilon>0$, the ground state is in SX phase with $\langle x\rangle>0$; when $\epsilon<0$, the ground state is in SP phase with $\langle p\rangle>0$.
+In the region $g_{r}<\left|\tilde{g}_{C}\right|$, both $\langle x\rangle$ and $\langle p\rangle$ are 0, which corresponds to the normal phase, as labeled by $N$.
+Outside the normal phase region, i.e., $g_{r}>\left|\tilde{g}_{C}\right|$, when $\epsilon>0$, the ground state is in the SX phase with $\langle x\rangle>0$; when $\epsilon<0$, the ground state is in the SP phase with $\langle p\rangle>0$.
 One can distinguish the phases by $\langle x\rangle$ and $\langle p\rangle$.
-Figure 6: (a) (Color online) (a-b) The first derivatives $\partial E_{G}/\partial\lambda$ and $\partial E_{G}/\partial\mu$ versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$.
-(b-c) The second derivatives $\partial^{2}E_{G}/\partial\lambda^{2}$ and $\partial^{2}E_{G}/\partial\mu^{2}$ versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$.
-Other parameters are $\omega_{0}=1,\eta=10^{2}$.
 To characterize the type of the QPTs, we calculate the derivatives of ground-state energy $E_{G}$ over the coupling strengths $g_{r}$ and $g_{cr}$.
@@
 We perform the squeezing transformation
-$S(r_{p})=\exp[\frac{r_{p}}{2}(a^{2}-a^{\dagger 2})]$ (34)
-on $\tilde{H}_{p}$ (p=N, SX, SP, SXPa and SXPb). The detailed form of $r_{p}$ is given in the Appendix B. Then the resulting Hamiltonian reads
+$S(r_{p})=\exp\left[\frac{r_{p}}{2}(a^{2}-a^{\dagger 2})\right]$ (35)
+on $\tilde{H}_{p}$ ($p$=N, SX, SP, SXPa, and SXPb). The detailed form of $r_{p}$ is given in the Appendix C. Then the resulting Hamiltonian reads
 $\tilde{H}(r_{p})=$ $S^{\dagger}(r_{p})\tilde{H}_{p}S(r_{p})$
@@
 $E_{G}^{\text{SP}}=$ $\frac{\omega_{\text{SP}}}{2}+\frac{\tilde{\omega}_{c}}{2}\frac{\zeta}{\zeta^{\prime 3}}-\frac{\tilde{\omega}_{0}}{4}(\zeta^{\prime 2}+\zeta^{\prime-2})-\frac{\tilde{\omega}_{c}}{2},$
-$E_{G}^{\text{SXPa}}=$ $\-\frac{\tilde{\omega}_{c}}{2}(1+\frac{4}{\mu^{2}})-\tilde{\omega}_{0}(\frac{\mu^{2}}{16}+\frac{1}{\mu^{2}}),$
-$E_{G}^{\text{SXPb}}=$ $\frac{\tilde{\omega}_{c}}{2}(1-\frac{4}{\lambda^{2}})-\tilde{\omega}_{0}(\frac{\lambda^{2}}{16}+\frac{1}{\lambda^{2}}).$
+$E_{G}^{\text{SXPa}}=$ $-\frac{\tilde{\omega}_{c}}{2}\left(1+\frac{4}{\mu^{2}}\right)-\tilde{\omega}_{0}\left(\frac{\mu^{2}}{16}+\frac{1}{\mu^{2}}\right),$
+$E_{G}^{\text{SXPb}}=$ $-\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{4}{\lambda^{2}}\right)-\tilde{\omega}_{0}\left(\frac{\lambda^{2}}{16}+\frac{1}{\lambda^{2}}\right).$
 We plot the dependence of the first and second derivatives of $E_{G}$ on $\lambda$ and $\mu$ versus $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ in Fig.
@@
 6(a) and 6(b), it is found that $\partial E_{G}/\partial\lambda$ ($\partial E_{G}/\partial\mu$) is discontinuous at $g_{r}=0\ (g_{cr}=0)\ \text{for}\ \left|g_{cr}/g_{C}\right|>2$ ($\left|g_{r}/g_{C}\right|>2$).
-This results denote the QPT from SX to SP is of first order.
+This result denotes the QPT from the SX phase to the SP phase is of first order.
 As shown in Figs.
@@
 Thus, the transition from the normal phase to superradiant phases is of second order.
-We also use the dashed and solid lines to indicate the first and second order QPT in Fig.
+We also use the dashed and solid lines to indicate the first- and second-order QPT in Fig.
 5, respectively.
@@
 7.
-The subscript “p” that labels the specific phase in $\omega_{p}$ is omitted.
-Figure 7: (Color online) The excitation energy $\omega$ versus the dimensionless modulation amplitude $\xi$. Parameters are $\eta=10^{2},\nu/\omega_{0}=0.68$ and $g/\omega_{0}=0.06$.
+The subscript $p$ that labels the specific phase in $\omega_{\text{p}}$ is omitted.
 The solid curve indicates that the system can enter different phases by adjusting the dimensionless modulation amplitude $\xi$.
-The QPT occurs at $\xi=1.10172$ and 2.59789, where $g_{r}/g_{C}=-1$, $g_{cr}/g_{C}=-1$.
-The ground state is in the normal phase when $1.101720<\xi<$ 1.10172 and $32.59789<\xi<$ 3, and the SX phase appears when $2.597891.10172<\xi<$ 2.59789.
-The system experiences the QPTs from the normal phase to SX phase at $\xi=1.10172$ and from SX phase to the normal phase at $\xi=2.59789$.
-To see the system how to cross the phase boundary under the chosen parameter in Fig.
+The QPT occurs at $\xi=1.102$ and $2.598$, where $g_{r}/g_{C}=-1$, $g_{cr}/g_{C}=-1$.
+The ground state is in the normal phase when $1.1020<\xi<$ 1.102 and $32.598<\xi<$ 3, and the SX phase appears when $2.5981.102<\xi<$ 2.598.
+The system experiences the QPTs from the normal phase to the SX phase at $\xi=1.102$ and from the SX phase to the normal phase at $\xi=2.598$.
+Figure 9: (Color online) The ratio $g_{A^{2}}/2\omega_{c}^{\prime}$ versus the coupling strength $g/\omega_{0}$ and the dimensionless factor $\chi$.
+Parameters are (a) $\omega_{0}/\omega_{c}=10^{2}$ and (b) $\omega_{0}/\omega_{c}=1$.
+To see the system how crosses the phase boundary under the chosen parameter in Fig.
 7, we plot the parameterized curves of [$g_{r}(\xi)/g_{C}$, $g_{cr}(\xi)/g_{C}$] versus the modulation amplitude $\xi$ in Fig.
 8(a).
-The parameterized curve evolves from the normal phase to SX phase then back to the normal phase as $\xi$ increases from 0 to 3.
+The parameterized curve evolves from the normal phase to the SX phase and then back to the normal phase as $\xi$ increases from 0 to 3.
 The system can be described by a standard QRM.
@@
 As shown in Fig.
-8(b), although the coupling strength is only in the strong-coupling regime with $g/\omega_{0}=0.06$, yet by the modulation, the QRM can reach the superradiant phase when $2.597891.10172<\xi<$ 2.59789.
-The normal phase appears when $1.101720<\xi<$ 1.10172 and $32.59789<\xi<$ 3.
+8(b), although the coupling strength is only in the strong-coupling regime with $g/\omega_{0}=0.06$, yet by the modulation, the QRM can reach the superradiant phase when $2.5981.102<\xi<$ 2.598.
+The normal phase appears when $1.1020<\xi<$ 1.102 and $32.598<\xi<$ 3.
 The results shown in Fig.
@@
 Thus, we can manipulate quantum phases on the QRM via the appropriate modulation with strong coupling.
-Figure 8: (Color online) (a) The parameterized curve of the couplings $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ with the modulation amplitude $\xi$ increasing from 0 to 3 as indicated by the arrows.
-(b) $\langle x\rangle$ versus the modulation amplitude $\xi$.
-The parameters are the same as Fig.
-7.
-We mention that, the anisotropic QRM can be realized with our scheme, as shown in Fig.
+We mention that the anisotropic QRM can be realized with our scheme, as shown in Fig.
 2(b).
@@
 Despite this, we still provide a possible scheme to observe the superradiant phase of few-body quantum systems in the strong-coupling regime.
-In this paper, we consider negligible cavity dissipation with $\kappa/\omega_{0}<10^{-4},\ \kappa/\omega_{c}^{\prime}<10^{-2}$.
-When the cavity dissipation is non-negligible, it will produce a shift to the critical coupling strength, which becomes $g_{C}^{\text{diss}}=\frac{1}{2}\sqrt{\frac{\tilde{\omega}_{0}}{\tilde{\omega}_{c}}(\tilde{\omega}_{c}^{2}+\kappa^{2})}$.
+In this paper, we consider a negligible cavity dissipation with $\kappa/\omega_{0}<10^{-4}$ and $\kappa/\omega_{c}^{\prime}<10^{-2}$ Johansson ; RMP ; Probing .
+When the cavity dissipation is non-negligible, it will produce a shift to the critical coupling strength, which becomes $g_{C}^{\text{diss}}=\frac{1}{2}\sqrt{\frac{\tilde{\omega}_{0}}{\tilde{\omega}_{c}}(\tilde{\omega}_{c}^{2}+\kappa^{2})}$ dispassive QRM .
 For a dissipation rate $\kappa=0.2\omega_{c}^{\prime}=0.002\omega_{0}$ and $\nu/\omega_{0}=0.68$, $\tilde{\omega}_{c}=0.01\omega_{0}$ and $g_{C}^{\text{diss}}\approx 1.02g_{C}$.
@@
 Hence, it is still effective for our approach to manipulate the steady-state superradiant phase transition of the dissipative QRM.
-VII two-photon $A^{2}$ terms
-In this section, we discuss the validity of omitting the two-photon $A^{2}$ terms. The two-photon $A^{2}$ terms given by Hamiltonian (5) can be written as
+VII Analysis concerning the two-photon $A^{2}$ term
+It is known that the existence of superradiant phase transition in the cavity-QED system is subject to a no-go theorem, where the squared term of the electromagnetic vector potential (i.e., the $A^{2}$ term) leads to the disappearance of the QPT in the Dicke model A2 term .
+This no-go theorem also applies to the QPT of the QRM Hwang2015Rabi .
+In the cavity-QED system, the coefficient $\chi$ in the $A^{2}$ term satisfies $\chi\geq 1$ due to Thomas-Reiche-Kuhn sum rule no-go theorm .
+Thus, the occurrence of the superradiant phase transition is still challenging in the cavity-QED system.
+However, in the circuit-QED system, $\chi$ could be smaller than 1.
+Then the superradiant phase transition is recovered with a shifted critical point caused by the $A^{2}$ term A2 term ; Hwang2015Rabi ; no-go theorm ; 1A2 ; 2A2 ; 3A2 .
+In this section, we discuss the validity of omitting the two-photon $A^{2}$ term. The two-photon $A^{2}$ term given by Hamiltonian (5) can be written as
 $H_{A^{2}}=2g_{A^{2}}a^{\dagger}a+g_{A^{2}}(a^{\dagger 2}+a^{2}).$ (38)
-The first term in $H_{A^{2}}$ has been absorbed into the Hamiltonian of the standard QRM by replacing the bosonic mode frequency $\omega_{c}$ with the modified frequency $\omega_{c}^{\prime}=\omega_{c}+2g_{A^{2}}$.
+The first term in $H_{A^{2}}$ is absorbed into the Hamiltonian of the standard QRM by replacing the bosonic mode frequency $\omega_{c}$ with the modified frequency $\omega_{c}^{\prime}=\omega_{c}+2g_{A^{2}}$.
 The second term can be safely neglected with $g\ll\omega_{0},\ g_{A^{2}}\ll 2\omega_{c}^{\prime}$.
@@
 9, we plot the ratio $g_{A^{2}}/2\omega_{c}^{\prime}$ versus the coupling strength $g$ and the dimensionless coefficient $\chi$.
-For the dispersion case $\eta=10^{2}$, as shown in Fig.
-9(a), to satisfy $g_{A^{2}}/2\omega_{c}^{\prime}=0.01$, as indicated by the dashed curve, $g$ and $\chi$ need to be ($g/\omega_{0}\sim 0.06,\ \chi\sim 0.0579$) or ($g/\omega_{0}\sim 0.0102,\ \chi\sim 2$).
-For the parameters we used in our discussions with $\eta=10^{2}$, $g/\omega_{0}=0.06$, and $\chi\leq 0.05$, this ratio is smaller than $1\%$, namely, $g_{A^{2}}/2\omega_{c}^{\prime}<0.01.$ Hence, we can neglect these terms in the dispersion limit.
-For the resonance case $\eta=1$, as shown in Fig.
+For the dispersion case $\omega_{0}/\omega_{c}=10^{2}$, as shown in Fig.
+9(a), to satisfy $g_{A^{2}}/2\omega_{c}^{\prime}=0.01$, as indicated by the dashed curve, $g$ and $\chi$ need to be ($g/\omega_{0}\sim 0.06,\ \chi\sim 0.058$) or ($g/\omega_{0}\sim 0.01,\ \chi\sim 2$).
+For the parameters we use in our discussions with $\omega_{0}/\omega_{c}=10^{2}$, $g/\omega_{0}=0.06$, and $\chi\leq 0.05$, this ratio is smaller than $1\%$, namely, $g_{A^{2}}/2\omega_{c}^{\prime}<0.01.$ Hence, we can neglect the $A^{2}$ term in the dispersion limit under the chosen parameters.
+For the resonance case $\omega_{0}/\omega_{c}=1$, as shown in Fig.
 9(b), to satisfy $g_{A^{2}}/2\omega_{c}^{\prime}=0.01$, we should have ($g/\omega_{0}\sim 0.06,\ \chi\sim 5.787$).
-Therefore, it is also reasonable to omit the $A^{2}$ terms for $g/\omega_{0}=0.06$ and $\chi<5$ for resonance case $\eta=1$.
-Thus we can omit the $A^{2}$ terms from the following discussions.
-Figure 9: (Color online) The ratio $g_{A^{2}}/2\omega_{c}^{\prime}$ versus the coupling strength $g/\omega_{0}$ and the dimensionless factor $\chi$. Parameters are (a) $\eta=10^{2}$ and (b) $\eta=1$.
-VIII conclusion and discussions
-By applying a qubit frequency modulation to a standard QRM, we generate an anisotropic QRM whose effective couplings for rotating and counter-rotating terms can be adjusted over a wide range.
-We calculate the ground-state and excitation energy of this model in terms of the modulation parameters.
-The phase diagram of the anisotropic QRM in the whole coupling parameter space and the type of the QPT are also obtained.
-We show that the superradiant phase of the QRM can be observed in the strong-coupling regime with the frequency modulation.
-Furthermore, the two-photon $A^{2}$ terms, which could prevent the observation of the QPT in QRM Hwang2015Rabi , can be appropriately neglected in our scheme.
-We provide an experimentally accessible scheme to control the QPT in QRM.
+Therefore, it is also reasonable to omit the $A^{2}$ term for $g/\omega_{0}=0.06$ and $\chi<5$ for the resonance case $\omega_{0}/\omega_{c}=1$.
+Thus we can omit the $A^{2}$ term by choosing appropriate parameters.
+VIII Discussions on the experimental implementation
+In this section, we present some discussions concerning the experimental implementation of our scheme.
+The key element for the experimental implementation of this scheme is to realize the QRM with the strong coupling between the qubit and bosonic field, as well as the qubit frequency modulation.
+Here, we present some detailed discussions concerning the realization of the two points.
+On the one hand, the QPT of the QRM in our scheme occurs when the coupling strength between the qubit and the field is only in the strong-coupling regime, for example $g/\omega_{0}=0.06$.
+This coupling strength can be realized in various experimental platforms 3A2 ; ZhengShiBiao ; trapped ions ; LiJian ; Probing ; D.
+Lv ; cqed .
+To present some analyses on the current experimental conditions in these systems, the related experimental parameters reported in the circuit-QED system ZhengShiBiao and the trapped-ion system trapped ions are listed in Table 2.
+The frequency ratio is $\eta\simeq 10$ and 25 for the circuit-QED system ZhengShiBiao and the trapped-ion system trapped ions , respectively, which provide the condition for the occurrence of the QPT in QRM.
+On the other hand, the qubit frequency modulation is also accessible in several experimental platforms.
+For a superconducting circuit system, the transition frequency of an Xmon qubit, which couples to a resonator, can be periodically modulated by an ac magnetic flux  ZhengShiBiao .
+In a trapped-ion system, we can use a linear Pauli trap to confine an ion to simulate the QRM.
+A pair of counter-propagating Raman laser beams can be used to modulate the transition frequency of the qubit formed by hyperfine states trapped ions .
+As reported in LiJian , the frequency of the qubit is $\omega_{0}/2\pi=2.62$ GHz and the modulation amplitude is $\xi\nu/2=2\pi\times 250$ MHz.
+One can vary the modulation frequency $\nu/2\pi$ from 0 to 500 MHz, leading to $0\leq\nu/\omega_{0}\leq 0.19$ and $\xi>1$.
+Thus, the modulation parameters reported in experiments are of the same order of magnitude as those used in our scheme.
+We list the parameters for simulations of our scheme in Table 3.
+All the parameters are achievable in experiments trapped ions ; LiJian ; ZhengShiBiao ; D.
+Lv .
+In addition, we consider the negligible dissipation of the field and qubit.
+This assumption can also be supported by current experimental technology, based on the fact that the dimensionless decay rates of the field and qubit could be smaller than $10^{-3}$ in the above experimental platforms Johansson ; RMP ; Probing .
+IX conclusion
+We proposed a scheme for generating an anisotropic QRM by applying a qubit frequency modulation to a standard QRM.
+Here the effective couplings for both the rotating and counter-rotating terms in the obtained anisotropic QRM can be adjusted over a wide range.
+We calculated the ground-state and excitation energy of this model in terms of the modulation parameters.
+The phase diagram of the anisotropic QRM in the entire coupling parameter space and the type of the QPT were also obtained.
+It was shown that the superradiant phase of the QRM can be observed in the strong-coupling regime with the frequency modulation.
+Furthermore, the two-photon $A^{2}$ term, which could prevent the observation of the QPT in QRM Hwang2015Rabi , can be appropriately neglected in our scheme.
+We also provided an experimentally accessible scheme to control the QPT in QRM.
 Our scheme is within current experimental technology trapped ions ; nuclear ; ZhengShiBiao ; 24.2 .
 This modulation scheme may also be applicable in other few-body quantum systems.
-Our work may pave the way for experimentally observing superradiant phase in other few-body quantum systems.
+Our work will pave the way for experimentally observing the superradiant phase in other few-body quantum systems.
 Acknowledgements.
-J.-F.H. is supported in part by the National Natural Science Foundation of China (Grant No. 12075083).
-Appendix A Operators for superradiant phase
+J.-F.H.
+is supported in part by the National Natural Science Foundation of China (Grants No.
+12075083 and No.12475016) and the Key Program of Xiangjiang-Laboratory in Hunan province, China (Grant No.
+XJ2302001).
+We would like to thank Jie-Qiao Liao for the helpful discussions during the reply to the referees’ reports.
+Appendix A Evaluation of the validity of the rotating-wave approximation
+To evaluate the validity of the rotating wave approximation used in Hamiltonian (9), we examine the time evolution of the fidelity $F(t)=\left|\left\langle\phi\left(t\right)\right|\left.\psi\left(t\right)\right\rangle\right|^{2}$ between the states $\left|\phi\left(t\right)\right\rangle$ and $\left|\psi\left(t\right)\right\rangle$, which are governed by the Hamiltonian (9) and Hamiltonian (12), respectively.
+Here, we choose the initial state $\left|\phi\left(0\right)\right\rangle=\left|\psi\left(0\right)\right\rangle=(1/\sqrt{2})\left(\left|g\right\rangle+\left|e\right\rangle\right)\left|\alpha\right\rangle$, where $\left|\alpha\right\rangle$ is a coherent state of the cavity field.
+Table 2: Experimental parameters reported in the circuit-QED system ZhengShiBiao and trapped-ion system trapped ions : the transition frequency $\omega_{0}$ of the two-level system, the frequency $\omega_{c}$ of the cavity field, the coupling strength $g$, and the ratio $g/\omega_{0}$.
+Reference Description $\omega_{0}/2\pi$ $\omega_{c}/2\pi$ $g/2\pi$ $g/\omega_{0}$
+ZhengShiBiao Circuit QED $3.42-10.25$ MHz $0.34-1.03$ MHz 0.81 MHz $0.079-0.237$
+trapped ions Trapped ion 50 kHz 2 kHz $0-7.1$ kHz $0-0.142$
+Table 3: Analysis of the used parameters in our scheme corresponding to the experimental parameters given in Table 2.
+Used parameters Circuit QED Trapped ion
+[…]
+[…]
+[…]
+[…]
+Dimensionless modulation amplitude $\xi$ $\xi=0-3$- --
+[…]
+We plot the fidelity $F(t)$ versus $t$ for $\nu/\omega_{0}=0.49,\ 0.68,\text{and}\ 1$.
+As shown in Fig.
+10, the fidelity can be higher than 0.9 when $\omega_{0}t/2\pi\leq 9$ for the used parameters.
+Specifically, for $\nu/\omega_{0}=0.68,$ which is used in our simulations, the fidelity is higher than 0.97 for $\omega_{0}t/2\pi\leq 10$.
+In particular, the fidelity can be further improved by choosing a suitably larger $\nu$.
+This result implies that the RWA is valid in our scheme under the chosen parameters.
+Appendix B Operators for superradiant phase
 The form of displacing operator $D_{k}(\alpha)$ ($k$=SX, SP, SXPa, SXPb) for each $k$ phase reads
 $\begin{array}[]{r@{}l}D_{\text{SX}}(\alpha)&{}=e{}^{-\alpha\frac{\partial}{\partial x}},\\ D_{\text{SP}}(\alpha)&{}=e^{-\alpha\frac{\partial}{\partial p}},\\ D_{\text{SXPa}}(\alpha)&{}=e^{-\sqrt{2}\alpha\frac{\partial}{\partial x}},\\ D_{\text{SXPb}}(\alpha)&{}=e^{-\sqrt{2}\alpha\frac{\partial}{\partial x}}.\end{array}$ (39)
-The new Pauli operators we introduced are
+The new Pauli operators we introduce are
 $\begin{array}[]{r@{}l}\tau_{x}^{\text{SX}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SX}}}\sigma_{x}-\frac{1}{\Omega_{\text{SX}}}\sqrt{2\tilde{\omega}_{0}\tilde{\omega}_{c}}\zeta\left|\alpha_{\text{SX}}\right|\sigma_{z},\\ \tau_{y}^{\text{SX}}&{}=\sigma_{y},\\ \tau_{z}^{\text{SX}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SX}}}\sigma_{z}+\frac{1}{\Omega_{\text{SX}}}\sqrt{2\tilde{\omega}_{0}\tilde{\omega}_{c}}\zeta\left|\alpha_{\text{SX}}\right|\sigma_{x}\end{array}$ (40)
-for SX phase,
+for the SX phase,
 $\begin{array}[]{r@{}l}\tau_{x}^{\text{SP}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SP}}}\sigma_{y}+\frac{1}{\Omega_{\text{SP}}}\sqrt{2\tilde{\omega}_{0}\tilde{\omega}_{c}}\zeta^{\prime}\left|\alpha_{\text{SP}}\right|\sigma_{z},\\ \tau_{y}^{\text{SP}}&{}=-\sigma_{x},\\ \tau_{z}^{\text{SP}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SP}}}\sigma_{z}-\frac{1}{\Omega_{\text{SP}}}\sqrt{2\tilde{\omega}_{0}\tilde{\omega}_{c}}\zeta^{\prime}\left|\alpha_{\text{SP}}\right|\sigma_{y}\end{array}$ (41)
-for SP phase,
+for the SP phase,
 $\begin{array}[]{r@{}l}\tau_{x}^{\text{SXPa}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPa}}}\sigma_{x}-\frac{1}{\Omega_{\text{SXPa}}}2g_{c}\left|\alpha_{\text{SXPa}}\right|\sigma_{z},\\ \tau_{y}^{\text{SXPa}}&{}=\sigma_{y},\\ \tau_{z}^{\text{SXPa}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPa}}}\sigma_{z}+\frac{1}{\Omega_{\text{SXPa}}}2g_{c}\left|\alpha_{\text{SXPa}}\right|\sigma_{x}\end{array}$ (42)
-for SXPa phase and
+for the SXPa phase, and
 $\begin{array}[]{r@{}l}\tau_{x}^{\text{SXPb}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPb}}}\sigma_{x}-\frac{1}{\Omega_{\text{SXPb}}}2g_{r}\left|\alpha_{\text{SXPb}}\right|\sigma_{z},\\ \tau_{y}^{\text{SXPb}}&{}=\sigma_{y},\\ \tau_{z}^{\text{SXPb}}&{}=\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPb}}}\sigma_{z}+\frac{1}{\Omega_{\text{SXPb}}}2g_{r}\left|\alpha_{\text{SXPb}}\right|\sigma_{x}\end{array}$ (43)
-for SXPb phase, respectively. The form of unitary operator $U_{k}$ (k=SX, SP, SXPa, SXPb) for k phase reads
+for the SXPb phase, respectively.
+The form of unitary operator $U_{\text{k}}$ ($k$=SX, SP, SXPa, SXPb) for $k$ phase reads
 $U_{\text{SX}}$ $=\exp[\frac{-i}{\sqrt{2}{\Omega_{\text{SX}}}}\sqrt{\tilde{\omega}_{0}\tilde{\omega}_{c}}(x\tau^{{}_{\text{SX}}}_{y}/\zeta+\zeta^{\prime}p\tau_{x}^{{}_{\text{SX}}})],$ (44)
 $U_{\text{SP}}$ $=\exp[\frac{i}{\sqrt{2}{\Omega_{\text{SP}}}}\sqrt{\tilde{\omega}_{0}\tilde{\omega}_{c}}(-\zeta x\tau_{x}^{{}_{\text{SP}}}+p\tau_{y}^{{}_{\text{SP}}}/\zeta^{\prime})],$
-$U_{\text{SXPa}}$ $=\exp[\frac{-ig_{c}}{{\sqrt{2}\Omega_{\text{SXPa}}}}(\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPa}}}\tau_{y}^{{}_{\text{SXPa}}}x-\tau_{x}^{{}_{\text{SXPa}}}p)],$
-$U_{\text{SXPb}}$ $=\exp[\frac{-ig_{r}}{{\sqrt{2}\Omega_{\text{SXPb}}}}(\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPb}}}\tau_{y}^{{}_{\text{SXPb}}}x+\tau_{x}^{{}_{\text{SXPb}}}p)].$
-Appendix B Squeezing parameters
+$U_{\text{SXPa}}$ $=\exp\left[\frac{-ig_{c}}{{\sqrt{2}\Omega_{\text{SXPa}}}}\left(\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPa}}}\tau_{y}^{{}_{\text{SXPa}}}x-\tau_{x}^{{}_{\text{SXPa}}}p\right)\right],$
+$U_{\text{SXPb}}$ $=\exp\left[\frac{-ig_{r}}{{\sqrt{2}\Omega_{\text{SXPb}}}}\left(\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPb}}}\tau_{y}^{{}_{\text{SXPb}}}x+\tau_{x}^{{}_{\text{SXPb}}}p\right)\right].$
+Figure 10: (Color online) Dynamics of the fidelity $F(t)$ for $\nu/\omega_{0}=0.49,\ 0.68,\text{and}\ 1$, respectively.
+The initial state of the system is $\left|\phi\left(0\right)\right\rangle=\left|\psi\left(0\right)\right\rangle=(1/\sqrt{2})\left(\left|g\right\rangle+\left|e\right\rangle\right)\left|\alpha\right\rangle$ with $\alpha=0.1$.
+We take the modulation amplitude $\xi=1.5$.
+Other parameters are the same as Fig.
+7.
+Appendix C Squeezing parameters
 The form of the squeezing parameter $r_{p}$ ($p$=N, SX, SP, SXPa, SXPb) is given by